### v1

# Learning Distances from Data with Normalizing Flows and Score Matching

Abstract
Density-based distances (DBDs) offer an elegant solution to the problem of metric learning. By defining a Riemannian metric which increases with decreasing probability density, shortest paths naturally follow the data manifold and points are clustered according to the modes of the data. We show that existing methods to estimate Fermat distances, a particular choice of DBD, suffer from poor convergence in both low and high dimensions due to i) inaccurate density estimates and ii) reliance on graph-based paths which are increasingly rough in high dimensions. To address these issues, we propose learning the densities using a normalizing flow, a generative model with tractable density estimation, and employing a smooth relaxation method using a score model initialized from a graph-based proposal. Additionally, we introduce a dimension-adapted Fermat distance that exhibits more intuitive behavior when scaled to high dimensions and offers better numerical properties. Our work paves the way for practical use of density-based distances, especially in high-dimensional spaces.
1 Introduction
Metric learning aims to determine a distance metric that accurately measures the similarity or dissimilarity between data points. Typically, data points are mapped into a space where simple metrics like Euclidean distance can be applied. Although this method is computationally efficient, it is limited by the configurations possible in Euclidean space, where arbitrary distance relationships between sets of points cannot be represented.
A more flexible but computationally expensive approach involves defining a Riemannian metric in the data space and solving for geodesic distances. This allows for arbitrary distance relationships, but it requires choosing an appropriate Riemannian metric. Fermat distances (Groisman et al., 2022), which are a type of density-based distance (Bousquet et al., 2003), offer an elegant solution: the metric should be conformal (a multiple of the identity matrix) and inversely proportional to the probability density. In this way, geodesics pass through high-density regions of the data, respecting any inherent manifold structure. This approach is consistent with Fermat’s principle of least time in optics, where the inverse density functions like a refractive index.
To address the computational challenge of computing geodesics with this metric, previous work has approximated trajectories using shortest paths in a graph of nearest neighbors (Bijral et al., 2012). The edge weights are approximated by raising Euclidean distances to an appropriate power, with theoretical analysis supporting the consistency of this approximation. However, in practice, we observe that convergence quickly plateaus, showing no significant improvement with increasing sample size.
By employing more accurate edge weights obtained through normalizing flows, we achieve much faster convergence. Additionally, previous approaches did not account for the limitations of graph-based methods in higher dimensions, where paths become increasingly rough due to the curse of dimensionality. To address this, we use a relaxation scheme to smooth trajectories and maintain satisfactory convergence. We find that training models via score matching is more effective for this purpose than using normalizing flows.
In summary, our contributions are threefold: First, we introduce the use of normalizing flows to obtain more accurate edge weights for estimating Fermat distances, significantly improving convergence rates. Second, we address the limitations of graph-based methods in high-dimensional spaces by implementing a smooth relaxation scheme using score matching, which better maintains the continuity and smoothness of the trajectories. Finally, we propose a dimension-adapted Fermat distance that scales intuitively with increasing dimensionality and exhibits superior numerical properties, paving the way for practical applications of density-based distances in complex, high-dimensional datasets. Through these innovations, we demonstrate the practicality and effectiveness of our approach, opening up a new possibility for metric learning using density-based distances in high-dimensional spaces.
Figure 1: Ground truth and estimated geodesics on a 2-dimensional spiral dataset. The green line shows a ground truth geodesic, solved via relaxation on the true density (Algorithm 1). The dashed orange line is an estimate for the geodesic, found as the shortest path in a nearest neighbor graph with edge weights calculated from a normalizing flow density (Algorithm 2).
2 Related work
Bousquet et al. (2003) introduces the idea of changing the geometry of the data space by use of a conformal density-based metric in order to get more informative distance functions. However they do not attempt to calculate geodesics or distances implied by this geometry, instead linking it to a modified SVM algorithm used for semi-supervised learning. Sajama and Orlitsky (2005) also uses a density-based metric for semi-supervised learning. They estimate densities using kernel density estimation and construct a graph of nearest neighbor distances from which they compute geodesics.
Bijral et al. (2012) continues this line of work by approximating the density as inversely proportional to a power of the Euclidean distance between close neighbors. Hwang et al. (2016) gives convergence guarantees for the power-weighted shortest path distances to geodesic distances under the density-dependent metric. Several other papers give similar results, and some practical implementations (Chu et al., 2020; Little et al., 2022; Groisman et al., 2022; Moscovich et al., 2017; Alamgir and Von Luxburg, 2012; Mckenzie and Damelin, 2019; Little et al., 2020). Terms used are power-weighted shortest path distances (PWSPD) (Little et al., 2022) and Fermat distances (Groisman et al., 2022; Trillos et al., 2023).
One crucial contribution of our work is to check that proposed methods actually converge in practice, regardless of theoretical guarantees. Previous work does not compare shortest paths to ground truth geodesics, even for a uniform distribution, where geodesics are simply straight lines. We address this major shortcoming by computing the ground truth geodesics for problems with known density and find that existing methods show very poor convergence.
3 Background
3.1 Riemannian geometry
Riemannian geometry offers powerful tools for analyzing curved spaces, which can be particularly useful for understanding and optimizing over non-Euclidean spaces such as manifolds. A fundamental object in this field is the metric tensor, where different choices of metric can lead to vastly different geometric structures, such as positive curvature (e.g., a sphere which closes in on itself) or negative curvature (e.g., hyperbolic space which expands exponentially quickly), as well as more complicated mixtures of the two.
More formally, a Riemannian manifold $\mathcal{M}$ is a smooth $n$-dimensional space that locally resembles $\mathbb{R}^{n}$. At each point $p$ on the manifold, there is an associated tangent space $\mathcal{T}_{p}$, which is a linear approximation of the manifold at $p$. The metric tensor $g$ is a smoothly varying positive-definite bilinear form on the tangent space. It defines an inner product in the tangent space, allowing us to measure lengths and angles.
[…]
The distance between two points on the manifold is defined as the minimal length of a curve connecting the points. The curves that minimize this length are known as geodesics, which are the generalization of straight lines in Euclidean space to curved spaces.
3.2 Fermat distances
Density-based distances (DBDs) provide a way to define distances between points in a distribution by utilizing the underlying probability density function. This approach can be particularly useful for tasks such as clustering and pathfinding in data analysis and machine learning. If designed well, a DBD can capture desirable properties such as:
  1. 1.
    Shortest paths between points should pass through supported regions of the data. If there is a region with no data between two points, the shortest path should deviate around it to pass through regions with more data.
  2. 2.
    Data should cluster according to its modes. If two data points belong to different modes with little data connecting them, the distance between them should be high.
One way to define a DBD that satisfies these properties is through Fermat distances. Named after Fermat’s principle of least time in optics, these distances model the shortest paths between points by considering the inverse probability density (or a monotonic function of it) as analogous to the refractive index.
In the language of Riemannian geometry, we define a metric tensor of the form
$g(u,v)=\frac{\langle u,v\rangle}{f(p)^{2}},$ (1)
[…]
The length of a smooth curve $\gamma(t)$ is therefore given by
[…]
3.2.1 Fermat distances which scale with dimension
Figure 2: Geodesics of the standard normal distribution, projected into 2d, taken between two orthogonal points at distance $\sqrt{D}$ from the origin. The paths are scaled by $\sqrt{D}$ to enable comparison. (Left) Unscaled metric (same temperature for all dimensions) leads to sharply curved geodesics in high dimensions. (Right) Scaled metric (temperature equal to dimension) leads to consistent geodesics in all dimensions (note that the lines are overlapping).
If $\beta=1$, this leads to unintuitive behavior in higher dimensions. Consider the standard normal distribution in $D$ dimensions. Due to rotational symmetry, the geodesic connecting $x_{1}$ and $x_{2}$ will lie in the plane spanned by these two points and the origin. Without loss of generality, suppose this plane aligns with the first two axis directions. The density restricted to this plane is the same as in the two-dimensional case, up to a constant factor. Therefore, geodesics in the $D$-dimensional standard normal distribution are the same as in the two-dimensional standard normal distribution, once we account for the rotation into the plane of $x_{1}$ and $x_{2}$.
However, despite this similarity, the typical distance from the origin of points sampled from the $D$-dimensional standard normal distribution is much larger, with a mean of $\sqrt{D}$. As a result, geodesics will start and end in very low-density regions relative to the two-dimensional standard normal. This means that most of the trajectory is spent moving towards and then outwards from the high-density region near the origin.
To address this issue, if we scale the temperature linearly with dimension ($\beta=1/D$), the results become consistent with the two-dimensional case (see Fig. 2). Scaling the temperature with dimension also increases numerical stability. This is because the effect of the probability density is reduced, leading to more Euclidean-like geodesics. This effect is clearly visible in Fig. 2, where the geodesics are much closer to straight lines in the scaled case.
3.3 Geodesic equation
From the metric, we can compute the geodesic equation. This is the second-order differential equation which all paths of minimum length (geodesics) must satisfy. In our case, we have the relatively simple solution for a trajectory $\gamma(t)$:
$\ddot{\gamma}-2\beta(s(\gamma)\cdot\dot{\gamma})\dot{\gamma}+\beta s(\gamma)\|\dot{\gamma}\|^{2}=0$ (3)
[…]
The geodesic equation has the property that the speed is constant with respect to the metric, meaning $\sqrt{g(\dot{\gamma},\dot{\gamma})}$ is constant. In other words, the geodesic distance traveled in a given time interval is the same at all points of the curve. If $\gamma$ passes from a high-probability region to a low-probability one, the distances in the low-probability region could be orders of magnitude larger than in the high-probability region. This makes numerically solving the equations challenging due to the vastly different Euclidean velocities along the trajectory.
It is convenient to reparameterize the equation such that the Euclidean velocity is constant instead, as this is much easier to handle numerically. This is especially useful for solving via relaxation, where having equal-length segments leads to much more stable behavior. Making this change results in a very similar equation (removing only the factor of 2 in the second term):
$\ddot{\varphi}-\beta(s(\varphi)\cdot\dot{\varphi})\dot{\varphi}+\beta s(\varphi)\|\dot{\varphi}\|^{2}=0$ (4)
See Section A.1 for derivations of these results.
4 Methods
4.1 Ground truth distances through relaxation
A relaxation scheme allows one to solve a differential equation constrained by boundary conditions by iteratively updating the solution such that it better obeys the differential equation. We divide the interval $[0,1]$ into equal time steps of size $h=1/n$ and write $\varphi_{i}=\varphi(ih)$. Initialize the values of the curve $\varphi$ at $t=ih$ for $i=0,\dots,n$, with the constraint that $\varphi_{0}$ and $\varphi_{n}$ are the chosen endpoints. The initial curve could be the straight line connecting the end points, or a more informative guess if additional information is available.
Equation 4 leads to the following relaxation scheme. Define
$v_{i}=\frac{\varphi_{i+1}-\varphi_{i-1}}{2}$ (5)
then update with
$\varphi^{\prime}_{i}=\frac{\varphi_{i+1}+\varphi_{i-1}}{2}+\frac{\beta}{2}\left(s(\varphi_{i})\|v_{i}\|^{2}-(s(\varphi_{i})\cdot v_{i})v_{i}\right)$ (6)
See Section A.2 for a derivation.
4.1.1 Implementation
In practice, some additional features are needed to ensure stable convergence of the relaxation scheme, especially when the score function is not particularly smooth. Firstly, we found that oscillatory features commonly arise. To address this, we use smoothing after the update, returning a value which is interpolated between the value at the node and the average of its neighbors. Secondly, for rougher score functions we use a step size between 0 and 1, where 0 corresponds to no update and 1 corresponds to the ordinary relaxation step. See Algorithm 1.
Algorithm 1 One step of relaxation. $\varphi=\{\varphi_{i}\}_{i=0}^{n}$ is the current state of the trajectory, $s$ is the score function, $\beta$ the inverse temperature, $\gamma\in[0,1)$ the smoothing parameter and $\alpha\in(0,1]$ the step size. This step should be iterated until convergence, which can be measured via the difference between $\varphi$ and $\varphi^{\prime}$.
function RelaxationStep($\varphi,s,\beta,\gamma,\alpha$)
     $v_{i}\leftarrow\frac{1}{2}(\varphi_{i+1}-\varphi_{i-1})$triangleright $i=1,\dots,n-1$
     $\varphi^{\prime}_{i}\leftarrow\frac{1}{2}(\varphi_{i+1}+\varphi_{i-1})+\frac{\beta}{2}\left(s(\varphi_{i})\|v_{i}\|^{2}-(s(\varphi_{i})\cdot v_{i})v_{i}\right)$
     $\varphi^{\prime}_{i}\leftarrow(1-\gamma)\varphi^{\prime}_{i}+\frac{\gamma}{2}(\varphi^{\prime}_{i+1}+\varphi^{\prime}_{i-1})$
     $\varphi^{\prime}_{i}\leftarrow\alpha\varphi^{\prime}_{i}+(1-\alpha)\varphi_{i}$
     return $\varphi^{\prime}$
end function
4.2 Weighted graph algorithm based on Euclidean distance
If we have a large number of samples from $p(x)$, we can approximate the density between close neighbors $x_{1}$ and $x_{2}$ as constant and hence the shortest path between them as a straight line. The density is roughly proportional to the inverse of the Euclidean distance, raised to the power $d$, where $d$ is the intrinsic dimension of the distribution (Bijral et al., 2012):
$p(\gamma)\propto\frac{1}{\|x_{1}-x_{2}\|^{d}}$ (7)
Hence the distance between $x_{1}$ and $x_{2}$ according to the metric is approximately proportional to a power of the Euclidean distance
$\operatorname{dist}(x_{1},x_{2})\propto\|x_{1}-x_{2}\|^{\beta d+1}$ (8)
Bijral et al. (2012) therefore proposes to approximate shortest paths between more distant points by constructing a $k$-nearest neighbors graph of the samples with edge weights given by $\|x_{1}-x_{2}\|^{\beta d+1}$ and applying Dijkstra’s algorithm.
Hwang et al. (2016) gives consistency guarantees for this estimate, proving that a scaled version of it converges to the ground truth distance with a rate of $\exp(-\theta_{0}n^{1/(3d+2)})$, with $n$ the sample size, $d$ the intrinsic dimension, and $\theta_{0}$ a positive constant. Groisman et al. (2022) additionally show that the shortest paths converge to geodesics in a large sample size limit.
While this approach is appealing in its simplicity, can be easily implemented using standard libraries, and comes with theoretical guarantees, the density estimate is too poor to provide accurate estimates of Fermat distances in practice. In the experimental section 5.1.1 we show that paths computed with this approach do not converge to the ground truth geodesics as sample size increases, even in very simple 2-dimensional distributions such as the standard normal.
We are unsure why we observe such a large deviation from the theoretical prediction. It is possible that the unknown constant $\theta_{0}$ is very small, leading to extremely slow convergence, despite the exponential function. We leave investigation into this point to future work.
4.3 Normalizing flows
Normalizing flows (Rezende and Mohamed, 2015; Kobyzev et al., 2020; Papamakarios et al., 2021) are a set of techniques to learn a probability density via a parameterized diffeomorphism $f_{\theta}$ which maps from data to latent space. The probability density is obtained through the change of variables formula:
$p_{\theta}(x)=\pi(f_{\theta}(x))\left|\det\left(\frac{\partial f_{\theta}}{\partial x}\right)\right|^{-1}$ (9)
where $\pi$ is a simple latent distribution. A common strategy to design normalizing flows is through coupling blocks (Dinh et al., 2014, 2016; Durkan et al., 2019), where $f_{\theta}$ is the composition of a series of blocks, each with a triangular Jacobian. This makes the determinant easy to calculate, and hence the above formula is tractable.
Normalizing flows are trained by minimizing the negative log-likelihood on a training set:
$\mathcal{L}=\mathbb{E}_{p_{\text{data}}(x)}[-\log p_{\theta}(x)]$ (10)
4.4 Density-weighted graph algorithm
In order to better estimate probability densities, we use a normalizing flow trained on the data and directly use the density predicted by the model. In order to approximate the geodesic distance between close neighbors, we assume that the geodesic connecting them is approximately a straight line, and estimate its length by a discrete approximation:
$\operatorname{dist}(x_{1},x_{2})\approx\sum_{i=1}^{S}\frac{\|y_{i}-y_{i-1}\|}{p_{\theta}(y_{i-1/2})}$ (11)
Here the $y_{i}$ are points which interpolate $x_{1}$ and $x_{2}$ with $y_{0}=x_{1}$ and $y_{S}=x_{2}$, and $y_{i-1/2}$ is the midpoint of $y_{i-1}$ and $y_{i}$.
This allows us to take a k-nearest neighbors graph of the data and replace the edge weights by the approximate distance. This results in a sparse graph, to which we can apply Dijkstra’s algorithm to estimate shortest paths and distances between any two points. See Algorithm 2.
Algorithm 2 Construction of a density-weighted graph which can subsequently be used to compute shortest paths and distances between points. $X$ is the data matrix, $p$ the probability density, $\beta$ the inverse temperature, $k$ the number of neighbors, and $S$ the number of segments with which to approximate the integral. The computations can also be performed in the log space, returning log edge weights, which is more numerically stable in higher dimensions.
function DensityGraph($X,p,\beta,k,S$)
     $\mathcal{G}\leftarrow\texttt{KNN}(X,k)$triangleright Construct $k$-nearest neighbor graph
     for $(l,m)\in\mathcal{G}$ $\triangleright$ Iterate over edges
         $y_{i}\leftarrow(1-\frac{i}{S})X_{l}+\frac{i}{S}X_{m}$$\triangleright$ For $i=0,\dots,S$
         $y_{i-1/2}\leftarrow\frac{1}{2}(y_{i-1}+y_{i})$$\triangleright$ For $i=1,\dots,S$
         $\mathcal{G}_{lm}\leftarrow\sum_{i=1}^{S}\|y_{i}-y_{i-1}\|/p(y_{i-1/2})^{\beta}$
     end for
     return $\mathcal{G}$
end function
4.5 Relaxation with a learned score model
The score function is the gradient of the log-likelihood with respect to spatial inputs, i.e., $s(x)=\nabla_{x}\log p(x)$. Note the difference from the definition commonly used in statistic, where the derivative is with respect to model parameters.
Score matching (Hyvärinen, 2005) is a family of methods to estimate this function from data, including variants such as sliced score matching (Song et al., 2020a), denoising score matching (Vincent, 2011) and diffusion models (Song et al., 2020b).
Sliced score matching
We can learn the score of $p(x)$ by minimizing the following objective (known as the score matching objective):
$\mathbb{E}_{p(x)}[\operatorname{tr}(\nabla_{x}s_{\theta}(x))+\tfrac{1}{2}\|s_{\theta}(x)\|^{2}]$ (12)
In high dimensions the trace term is expensive to evaluate exactly, so sliced score matching uses a stochastic approximation (the Hutchinson trace estimator (Hutchinson, 1989)):
$\mathbb{E}_{p(x)p(v)}[v^{T}\nabla_{x}s_{\theta}(x)v+\tfrac{1}{2}\|s_{\theta}(x)\|^{2}]$ (13)
where $v$ is sampled from an appropriate distribution, typically standard normal.
Since only the score is needed for the relaxation algorithm described in Section 4.1, we can use a trained score model to perform relaxation to approximate geodesics. In practice, we initialize the path from a graph-based method to speed up convergence and avoid local minima.
5 Experiments
Performance metric
After solving for a shortest path $\varphi$ between two points $x_{1}$ and $x_{2}$, we compute its length $L(\varphi)$ by using the ground truth density function, then compare this value to the ground truth distance between the two points:
$M(\varphi)=\log\frac{L(\varphi)}{\operatorname{dist}(x_{1},x_{2})}$ (14)
Provided we have access to the ground truth density, this performance metric can be calculated independently of how $\varphi$ is obtained, and has a value of zero only if $\varphi$ is the ground truth geodesic. The metric will otherwise take positive values.
5.1 Scaling the dataset
Figure 3: The length of shortest paths in power-weighted distance graphs do not converge to the ground truth with increasing sample size (green). In contrast, learned and ground-truth density weighted graphs (orange and green) do converge, at similar rates.
In this section we perform experiments on a set of 5 two-dimensional datasets for which we know the ground truth density (see Fig. 3). Details on how the datasets were created can be found in Appendix B.
5.1.1 Power-weighted graph paths do not converge with increasing sample size
We first estimate paths using the power-weighted graph method (Bijral et al., 2012), applied to a k-nearest neighbor graph. We find that in all 5 datasets, increasing the sample size leads to saturating performance (see Fig. 3, green lines). In Fig. 6 in the appendix we show that increasing the number of neighbors does not improve performance.
5.1.2 Poor convergence is due to poor density estimation
In order to investigate the root cause of the poor convergence, we designed other graph methods which use the nearest neighbor estimator for the density, but do not use power-weighted graph edges. We use 4 variants for the density estimate (see Section C.1). We find that all 4 have very similar performance to the power-weighted graph, whereas the equivalent estimators using the ground truth density quickly converge. See Section C.1 for plots supporting this claim.
5.1.3 Learning the density leads to convergent paths
Since the problem lies in poor density estimation, we hope to fix the problem by better density estimation. To this end, we use a normalizing flow to learn the density from samples, then use Algorithm 2 to construct a graph over which we compute shortest paths. We find that the paths obtained converge at close to the same rate as paths obtained by an equivalent method using the ground truth densities (Fig. 3 orange and blue lines). Full experimental details can be found in Appendix B.
5.2 Scaling the dimension
Figure 4: Mean log error with a sample size of 200,000. Graph-based paths (the first three methods) converge very slowly in higher dimensions with increasingly worse performance as dimension increases. When using the scaled metric ($\beta=1/D$), convergence improves over the unscaled metric, but all graph-based methods perform similarly in higher dimensions. Doing relaxation with a learned score function (red line) leads to much faster convergence, avoiding the curse of dimensionality.
Machine learning datasets often lie in high-dimensional spaces, so it is crucial that any method applied to them scales well, avoiding the curse of dimensionality as much as possible. We test this by taking one of our two-dimensional datasets, namely the standard normal, and scaling it up to 25 dimensions. We perform experiments both for the scaled and unscaled version of the metric.
5.2.1 Graph-based paths do not converge well in higher dimensions
Our first finding is that all graph-based methods show declining performance as the dimension increases, for both unscaled and scaled metrics (Fig. 4). This can be straightforwardly explained by the curse of dimensionality: with increasing dimension but fixed sample size, it is increasingly unlikely that points in the dataset lie close to the true trajectory. The shortest path in the graph becomes increasingly rough in contrast to the smooth ground truth trajectory.
5.2.2 Relaxation using a learned score function improves convergence rates
In order to obtain smoother paths, we turn to the same relaxation algorithm we use to solve for ground truth geodesics. Instead of using the ground truth score function, we use a learned score function. An obvious choice would be to use automatic differentiation to get the score from a trained normalizing flow, but we find the resulting scores are too noisy to be useful, inhibiting reliable convergence of the relaxation scheme. Instead, we learn a score model with sliced score matching, which results in a smoother function suitable for performing relaxation. Full experimental details can be found in Appendix B.
We find that this fixes the convergence problems, both for the unscaled and scaled metrics, though convergence is more reliable in the scaled case (Fig. 4). Please see Fig. 8 and Fig. 9 in the appendix for more detailed evidence.
6 Conclusion
In conclusion, our study focuses on understanding the convergence properties of various algorithms to ground-truth geodesics by employing toy distributions with known densities. We investigate the standard normal distribution in dimensions higher than 2, highlighting the challenges faced by graph-based methods even in this simple scenario. Our findings demonstrate significant improvements through the use of normalizing flows and score models, yet further research is needed to address more complex distributions and real-world data where the underlying density is unknown. Future work should aim to unify the roles of normalizing flows and score models into a single cohesive model, enhancing the method’s applicability and efficiency. Additionally, a deeper theoretical analysis is required to understand why our proposed methods converge effectively, while previous methods with consistency guarantees do not. This work lays the groundwork for practical applications of density-based distances, setting the stage for more robust and scalable solutions in metric learning.
Appendix A Derivations
A.1 Geodesic equation
The geodesic equation which characterizes paths of zero acceleration $\gamma$ is
$\ddot{\gamma}^{i}+\Gamma^{i}_{jk}\dot{\gamma}^{j}\dot{\gamma}^{k}=0$ (15)
where $\Gamma^{i}_{jk}$ are the Christoffel symbols of the second kind
$\Gamma^{i}_{jk}=\frac{1}{2}g^{il}\left(\frac{\partial g_{lj}}{\partial x^{k}}+\frac{\partial g_{lk}}{\partial x^{j}}-\frac{\partial g_{jk}}{\partial x^{l}}\right)$ (16)
and $g^{il}$ is the inverse of $g$, i.e. such that $g^{il}g_{lj}=\delta^{i}_{j}$.
Our metric has the form $g_{ij}=\lambda^{2}\delta_{ij}$, meaning that
$\Gamma^{i}_{jk}$ $=\frac{1}{2\lambda^{2}}\delta^{il}\left(\frac{\partial\lambda^{2}}{\partial x^{k}}\delta_{lj}+\frac{\partial\lambda^{2}}{\partial x^{j}}\delta_{lk}-\frac{\partial\lambda^{2}}{\partial x^{l}}\delta_{jk}\right)$ (17)
$=\delta^{il}\left(\frac{\partial\log\lambda}{\partial x^{k}}\delta_{lj}+\frac{\partial\log\lambda}{\partial x^{j}}\delta_{lk}-\frac{\partial\log\lambda}{\partial x^{l}}\delta_{jk}\right)$ (18)
$=\frac{\partial\log\lambda}{\partial x^{k}}\delta^{i}_{j}+\frac{\partial\log\lambda}{\partial x^{j}}\delta^{i}_{k}-\frac{\partial\log\lambda}{\partial x^{l}}\delta^{il}\delta_{jk}$ (19)
In this case $\lambda=p^{-\beta}$ and hence $\frac{\partial\log\lambda}{\partial x}=-\beta\frac{\partial\log p}{\partial x}=-\beta s$, where $s=\frac{\partial\log p}{\partial x}$ is the score function of the probability density. As a result
$\Gamma^{i}_{jk}=-\beta s_{k}\delta^{i}_{j}-\beta s_{j}\delta^{i}_{k}+\beta s_{i}\delta^{il}\delta_{jk}$ (20)
Substituting into the geodesic equation leads to
$\ddot{\gamma}^{i}-\beta s_{k}\dot{\gamma}^{i}\dot{\gamma}^{k}-\beta s_{j}\dot{\gamma}^{j}\dot{\gamma}^{i}+\beta s_{l}\delta^{il}\delta_{jk}\dot{\gamma}^{j}\dot{\gamma}^{j}=0$ (21)
which can be written in vector notation as
$\ddot{\gamma}-2\beta(s(\gamma)\cdot\dot{\gamma})\dot{\gamma}+\beta s(\gamma)\|\dot{\gamma}\|^{2}=0$ (22)
Let $f:[0,1]\rightarrow[0,1]$ be a strictly increasing diffeomorphism. Given a curve $\gamma:[0,1]\rightarrow\mathbb{R}^{n}$, consider the curve $\varphi:[0,1]\rightarrow\mathbb{R}^{n}$ defined such that $\varphi(u)=\gamma(f^{-1}(u))$ or equivalently $\varphi(f(t))=\gamma(t)$. The derivatives of the two curves are related by
$\dot{\varphi}\dot{f}=\dot{\gamma}$ (23)
and
[…]
where derivatives of $\varphi$ are with respect to $u=f(t)$. By substituting these expressions into the geodesic equation, we arrive at the equivalent equation for $\varphi$:
$\ddot{\varphi}\dot{f}^{2}+\dot{\varphi}\ddot{f}-2\beta(s(\gamma)\cdot\dot{\varphi}\dot{f})\dot{\varphi}\dot{f}+\beta s(\gamma)\|\dot{\varphi}\dot{f}\|^{2}=0$ (25)
We would like $\varphi$ to have a constant Euclidean speed (the magnitude but not necessarily the direction of the velocity is constant), meaning
$\frac{d}{du}\left(\frac{1}{2}\|\dot{\varphi}\|^{2}\right)=\dot{\varphi}\cdot\ddot{\varphi}=0$ (26)
Multiplying the geodesic equation by $\dot{\varphi}$, this is equivalent to requiring
$\|\dot{\varphi}\|^{2}\ddot{f}-2\beta(s(\gamma)\cdot\dot{\varphi})\|\dot{\varphi}\|^{2}\dot{f}^{2}+\beta(s(\gamma)\cdot\dot{\varphi})\|\dot{\varphi}\|^{2}\dot{f}^{2}=0$ (27)
implying that
$\ddot{f}=\beta(s(\gamma)\cdot\dot{\varphi})\dot{f}^{2}$ (28)
Resubstituting into the geodesic equation and dividing by $\dot{f}^{2}$ leads to the result:
$\ddot{\varphi}-\beta(s(\varphi)\cdot\dot{\varphi})\dot{\varphi}+\beta s(\varphi)\|\dot{\varphi}\|^{2}=0$ (29)
A.2 Relaxation scheme
The central finite difference approximations for the first and second derivatives are
$\dot{\varphi}_{i}\approx\frac{\varphi_{i+1}-\varphi_{i-1}}{2h}$ (30)
and
$\ddot{\varphi}_{i}\approx\frac{\varphi_{i+1}-2\varphi_{i}+\varphi_{i-1}}{h^{2}}$ (31)
By substituting this into the differential equation we have
$\frac{\varphi_{i+1}-2\varphi_{i}+\varphi_{i-1}}{h^{2}}-\beta(s(\varphi_{i})\cdot\dot{\varphi}_{i})\dot{\varphi}_{i}+\beta s(\varphi_{i})\|\dot{\varphi}_{i}\|^{2}\approx 0$ (32)
where $\dot{\varphi}_{i}$ is estimated using finite differences. We can rearrange this to
$\varphi_{i}\approx\frac{\varphi_{i+1}+\varphi_{i-1}}{2}+\frac{h^{2}\beta}{2}\left(s(\varphi_{i})\|\dot{\varphi}_{i}\|^{2}-(s(\varphi_{i})\cdot\dot{\varphi}_{i})\dot{\varphi}_{i}\right)$ (33)
and use the equation as an update rule, updating each position of the curve except the endpoints at each iteration.
Dividing by very small $h$ could lead to numerical instability. We can avoid dividing and multiplying by $h$ by the following update rule. First define
$v_{i}=\frac{\varphi_{i+1}-\varphi_{i-1}}{2}$ (34)
then update with
$\varphi_{i}=\frac{\varphi_{i+1}+\varphi_{i-1}}{2}+\frac{\beta}{2}\left(s(\varphi_{i})\|v_{i}\|^{2}-(s(\varphi_{i})\cdot v_{i})v_{i}\right)$ (35)
Appendix B Experimental Details
Figure 5: The 5 two-dimensional datasets.
Software libraries
We used the numpy [Harris et al., 2020] and scipy [Virtanen et al., 2020] Python libraries for basic numerical and graph operations and matplotlib [Hunter, 2007] for plotting. We used scikit-learn [Pedregosa et al., 2011] to compute nearest neighbor graphs and pytorch [Paszke et al., 2019] for training neural networks. We used FrEIA [Ardizzone et al., 2018-2022] for constructing normalizing flows.
Two-dimensional datasets
We use the following two-dimensional datasets. All are implemented as Gaussian mixture models. The first two are explicitly Gaussian, the final three are GMMs with 50 components fitted using scikit-learn to data generated by adding noise to certain geometric structures.
  • Standard normal
  • Gaussian mixture with 3 components
  • Circle
  • 1 spiral
  • 2 spirals
Normalizing flow training
We used the FrEIA library [Ardizzone et al., 2018-2022] to design normalizing flows using spline coupling blocks [Durkan et al., 2019]. We used the following hyperparameters:
  • Blocks: 12 for two-dimensional data, 5 for standard normal in $D$ dimensions
  • Spline bins: 10
  • Domain clamping in the splines: 5
  • ActNorm between blocks
  • Noise of $10^{-5}$ added to the training data
  • LR = $5\times 10^{-4}$
  • Weight decay = $10^{-6}$
  • Fully-connected coupling block subnets with 3 hidden layers of 64 dimensions each, with ReLU activation and BatchNorm
  • Training iterations (as a function of training set size $n$): $2500\times 2^{n//1000}$
  • Batch size = 256 if $n>400$, else $\lfloor n/3\rfloor$
  • Adam optimizer
Score model training
We use fully connected networks to predict the score, with 6 hiddens layer with the width depending on the dimension of the input. We use the following hyperparameters:
  • Hidden layer width (as a function of input dimension $D$)
    • $3\leq D\leq 5$: 128
    • $6\leq D\leq 8$: 170
    • $9\leq D\leq 25$: 200
  • LR = $10^{-3}$
  • Weight decay = $10^{-6}$
  • Softplus activation function
  • Training iterations (as a function of training set size $n$): $2500\times 2^{n//1000}$
  • Batch size = 256 if $n>400$, else $\lfloor n/3\rfloor$
  • Adam optimizer
In all cases, the chosen model checkpoint was that which achieved the lowest training loss.
Compute resources
For the two-dimensional datasets, we performed all experiments on a machine with a NVIDIA GeForce RTX 2070 GPU and 32 GB of RAM. It took approximately 24 hours to run all experiments.
For the higher dimensional standard normal distributions, we performed all experiments on a machine with two GPUs: a NVIDIA GeForce RTX 2070 and a NVIDIA GeForce RTX 2080 Ti Rev. A. The machine has 32 GB of RAM. It took approximately 24 hours to run all experiments.
In addition we did some preliminary experiments on the first machine, totalling approximately 48 hours of compute time.
Appendix C Additional experimental results
Figure 6: Increasing the number of nearest neighbors does not significantly improve the convergence of the power-weighted graph method on two-dimensional datasets.
C.1 Convergence of alternative graph edge weightings
Figure 7: Alternative edge weightings based on the nearest neighbor density estimator do not converge. The same weightings based on the ground truth density do converge. Experiments performend on the GMM two-dimensional dataset.
We try out other edge weights, as described in the main text. Given the nearest-neighbor density estimates $\hat{p}(X_{l})\propto(\min_{m}\|X_{l}-X_{m}\|)^{-d}$ for data points $X_{l}$, we approximate the graph edges by $\mathcal{G}_{lm}=\|X_{l}-X_{m}\|/\tilde{p}((X_{l}+X_{m})/2)$ where $\tilde{p}$ is a density estimate for the midpoint. We use 4 variants:
  1. 1.
    Inverse of mean: $\tilde{p}((X_{l}+X_{m})/2)=(\hat{p}(X_{l})+\hat{p}(X_{m}))/2$
  2. 2.
    Mean of inverse: $21/\tilde{p}((X_{l}+X_{m})/2)=(1/\hat{p}(X_{l})+1/\hat{p}(X_{m}))/$ 2
  3. 3.
    Max: $\tilde{p}((X_{l}+X_{m})/2)=\max(\hat{p}(X_{l}),\hat{p}(X_{m}))$
  4. 4.
    Min: $\tilde{p}((X_{l}+X_{m})/2)=\min(\hat{p}(X_{l}),\hat{p}(X_{m}))$
We find that all 4 have very similar performance to the power-weighted graph, whereas the equivalent estimators using the ground truth density quickly converge. See Fig. 7.
Figure 8: Detailed plot of convergence behavior in higher dimensions, unscaled metric. We see that all graph-based methods plateau in performance, whereas the relaxation-based method continues to show good convergence.
Figure 9: Detailed plot of convergence behavior in higher dimensions, scaled metric. We see that all graph-based methods plateau in performance, whereas the relaxation-based method continues to show good convergence.

### v2

@@
 Abstract
-Density-based distances (DBDs) offer an elegant solution to the problem of metric learning.
-By defining a Riemannian metric which increases with decreasing probability density, shortest paths naturally follow the data manifold and points are clustered according to the modes of the data.
-We show that existing methods to estimate Fermat distances, a particular choice of DBD, suffer from poor convergence in both low and high dimensions due to i) inaccurate density estimates and ii) reliance on graph-based paths which are increasingly rough in high dimensions.
-To address these issues, we propose learning the densities using a normalizing flow, a generative model with tractable density estimation, and employing a smooth relaxation method using a score model initialized from a graph-based proposal.
-Additionally, we introduce a dimension-adapted Fermat distance that exhibits more intuitive behavior when scaled to high dimensions and offers better numerical properties.
-Our work paves the way for practical use of density-based distances, especially in high-dimensional spaces.
+Density-based distances (DBDs) provide a principled approach to metric learning by defining distances in terms of the underlying data distribution.
+By employing a Riemannian metric that increases in regions of low probability density, shortest paths naturally follow the data manifold.
+Fermat distances, a specific type of DBD, have attractive properties, but existing estimators based on nearest neighbor graphs suffer from poor convergence due to inaccurate density estimates.
+Moreover, graph-based methods scale poorly to high dimensions, as the proposed geodesics are often insufficiently smooth.
+We address these challenges in two key ways.
+First, we learn densities using normalizing flows.
+Second, we refine geodesics through relaxation, guided by a learned score model.
+Additionally, we introduce a dimension-adapted Fermat distance that scales intuitively to high dimensions and improves numerical stability.
+Our work paves the way for the practical use of density-based distances, especially in high-dimensional spaces.
+Machine Learning, ICML
 1 Introduction
-Metric learning aims to determine a distance metric that accurately measures the similarity or dissimilarity between data points.
-Typically, data points are mapped into a space where simple metrics like Euclidean distance can be applied.
-Although this method is computationally efficient, it is limited by the configurations possible in Euclidean space, where arbitrary distance relationships between sets of points cannot be represented.
+Figure 1: Our method learns distance metrics that respect the underlying data distribution.
+Here, we show shortest paths between two points in a complex Gaussian mixture distribution.
+The ground truth geodesic (dashed red) and our graph-based approximation (solid blue) both naturally follow high-density regions where data samples (green dots) are concentrated, rather than taking a direct Euclidean path.
+This behavior is desirable for many machine learning tasks, as it captures the intrinsic structure of the data manifold.
+Metric learning is fundamental to many machine learning tasks, from clustering to classification.
+The goal is to determine a distance metric that accurately measures the similarity or dissimilarity between data points.
+While traditional approaches map data points into spaces where Euclidean distance can be directly applied, these methods are inherently limited by the geometric constraints of Euclidean space, making it impossible to represent arbitrary distance relationships between sets of points.
 A more flexible but computationally expensive approach involves defining a Riemannian metric in the data space and solving for geodesic distances.
@@
 Fermat distances (Groisman et al., 2022), which are a type of density-based distance (Bousquet et al., 2003), offer an elegant solution: the metric should be conformal (a multiple of the identity matrix) and inversely proportional to the probability density.
-In this way, geodesics pass through high-density regions of the data, respecting any inherent manifold structure.
+Intuitively, this means that distances are stretched in regions where data is sparse and compressed where data is dense.
+In this way, geodesics pass through high-density regions of the data, respecting any inherent manifold structure (see Figure 1).
 This approach is consistent with Fermat’s principle of least time in optics, where the inverse density functions like a refractive index.
-To address the computational challenge of computing geodesics with this metric, previous work has approximated trajectories using shortest paths in a graph of nearest neighbors (Bijral et al., 2012).
-The edge weights are approximated by raising Euclidean distances to an appropriate power, with theoretical analysis supporting the consistency of this approximation.
-However, in practice, we observe that convergence quickly plateaus, showing no significant improvement with increasing sample size.
-By employing more accurate edge weights obtained through normalizing flows, we achieve much faster convergence.
-Additionally, previous approaches did not account for the limitations of graph-based methods in higher dimensions, where paths become increasingly rough due to the curse of dimensionality.
-To address this, we use a relaxation scheme to smooth trajectories and maintain satisfactory convergence.
-We find that training models via score matching is more effective for this purpose than using normalizing flows.
-In summary, our contributions are threefold: First, we introduce the use of normalizing flows to obtain more accurate edge weights for estimating Fermat distances, significantly improving convergence rates.
-Second, we address the limitations of graph-based methods in high-dimensional spaces by implementing a smooth relaxation scheme using score matching, which better maintains the continuity and smoothness of the trajectories.
-Finally, we propose a dimension-adapted Fermat distance that scales intuitively with increasing dimensionality and exhibits superior numerical properties, paving the way for practical applications of density-based distances in complex, high-dimensional datasets.
-Through these innovations, we demonstrate the practicality and effectiveness of our approach, opening up a new possibility for metric learning using density-based distances in high-dimensional spaces.
-Figure 1: Ground truth and estimated geodesics on a 2-dimensional spiral dataset.
-The green line shows a ground truth geodesic, solved via relaxation on the true density (Algorithm 1).
-The dashed orange line is an estimate for the geodesic, found as the shortest path in a nearest neighbor graph with edge weights calculated from a normalizing flow density (Algorithm 2).
-2 Related work
-Bousquet et al.
-(2003) introduces the idea of changing the geometry of the data space by use of a conformal density-based metric in order to get more informative distance functions.
-However they do not attempt to calculate geodesics or distances implied by this geometry, instead linking it to a modified SVM algorithm used for semi-supervised learning.
-Sajama and Orlitsky (2005) also uses a density-based metric for semi-supervised learning.
-They estimate densities using kernel density estimation and construct a graph of nearest neighbor distances from which they compute geodesics.
-Bijral et al.
-(2012) continues this line of work by approximating the density as inversely proportional to a power of the Euclidean distance between close neighbors.
-Hwang et al.
-(2016) gives convergence guarantees for the power-weighted shortest path distances to geodesic distances under the density-dependent metric.
-Several other papers give similar results, and some practical implementations (Chu et al., 2020; Little et al., 2022; Groisman et al., 2022; Moscovich et al., 2017; Alamgir and Von Luxburg, 2012; Mckenzie and Damelin, 2019; Little et al., 2020).
-Terms used are power-weighted shortest path distances (PWSPD) (Little et al., 2022) and Fermat distances (Groisman et al., 2022; Trillos et al., 2023).
-One crucial contribution of our work is to check that proposed methods actually converge in practice, regardless of theoretical guarantees.
-Previous work does not compare shortest paths to ground truth geodesics, even for a uniform distribution, where geodesics are simply straight lines.
-We address this major shortcoming by computing the ground truth geodesics for problems with known density and find that existing methods show very poor convergence.
+This density-based approach to metric learning has evolved significantly over the past two decades.
+The foundational work of Bousquet et al.
+(2003) introduced the concept of conformal density-based metrics, though their focus was on semi-supervised learning rather than computing geodesics.
+Sajama & Orlitsky (2005) took the first steps toward practical implementation by using kernel density estimation and nearest-neighbor graphs for geodesic computation.
+A breakthrough came from Bijral et al.
+(2011), who approximated the density as inversely proportional to a power of the Euclidean distance between close neighbors.
+This elegant approximation sparked numerous theoretical extensions and practical implementations (Chu et al., 2020; Little et al., 2022; Groisman et al., 2022; Moscovich et al., 2017; Alamgir & Von Luxburg, 2012; Mckenzie & Damelin, 2019; Little et al., 2020), becoming known as power-weighted shortest path distances (PWSPD) (Little et al., 2022) or Fermat distances (Groisman et al., 2022; Trillos et al., 2023).
+While Hwang et al.
+(2016) provided theoretical convergence guarantees for these methods, notably absent from all previous work is any comparison to ground truth distances, even in simple cases where they can be computed exactly.
+Our empirical analysis reveals that existing approaches exhibit poor convergence in practice, highlighting a significant gap between theoretical guarantees and practical performance.
+Other notable works which use density-based metrics to inform the learning of data manifolds include metric flow matching (Kapusniak et al., 2024) and the Riemannian autoencoder (Diepeveen et al., 2024), which uses a metric defined in terms of score functions.
+To establish ground truth geodesics and distances, we develop a numerically stable relaxation method that directly solves the geodesic differential equations.
+By carefully handling the reparameterization of paths to maintain constant Euclidean velocity and employing appropriate numerical optimization techniques, we can compute accurate geodesics even in challenging cases with large density variations.
+This enables us, for the first time, to quantitatively evaluate the accuracy of different approximation methods by comparing them to true geodesics in distributions with known densities, such as mixtures of Gaussians.
+With this ground truth framework in place, we develop two key improvements to existing methods.
+First, by employing more accurate edge weights obtained through normalizing flows, we achieve much faster convergence in graph-based approximations.
+Second, we address the limitations of these graph-based methods in higher dimensions, where paths become increasingly rough due to the curse of dimensionality, by using a relaxation scheme to smooth trajectories.
+Interestingly, we find that training models via score matching is more effective than using normalizing flows for this purpose, revealing a trade-off between the accuracy of learned log densities and their gradients (scores).
+In summary, our contributions are fourfold: First, we introduce a numerically stable relaxation method to compute ground truth geodesics and distances.
+Second, we introduce the use of normalizing flows to obtain more accurate edge weights for estimating Fermat distances, achieving significantly faster convergence rates than previous methods.
+Third, we overcome the limitations of graph-based methods in high-dimensional spaces through a novel smooth relaxation scheme using score matching.
+Finally, we propose a Fermat distance that naturally scales with dimensionality, making density-based distances practical for real-world, high-dimensional applications.
+Our approach bridges the gap between theoretical elegance and practical applicability in metric learning.
 2 Background
 2.1 Riemannian Geometry
-Riemannian geometry offers powerful tools for analyzing curved spaces, which can be particularly useful for understanding and optimizing over non-Euclidean spaces such as manifolds.
-A fundamental object in this field is the metric tensor, where different choices of metric can lead to vastly different geometric structures, such as positive curvature (e.g., a sphere which closes in on itself) or negative curvature (e.g., hyperbolic space which expands exponentially quickly), as well as more complicated mixtures of the two.
+Riemannian geometry provides the mathematical foundation for understanding curved spaces and measuring distances within them.
+While Euclidean geometry suffices for flat spaces, many real-world datasets lie on curved manifolds where Euclidean distances can be misleading.
+Riemannian geometry offers powerful tools for analyzing such spaces through the metric tensor, which can encode different geometric structures ranging from positive curvature (like a sphere) to negative curvature (like hyperbolic space), or more complex combinations of both.
 More formally, a Riemannian manifold $\mathcal{M}$ is a smooth $n$-dimensional space that locally resembles $\mathbb{R}^{n}$.
-At each point $p$ on the manifold, there is an associated tangent space $\mathcal{T}_{p}$, which is a linear approximation of the manifold at $p$.
-The metric tensor $g$ is a smoothly varying positive-definite bilinear form on the tangent space.
+In this work, we make the simplifying assumption that $\mathcal{M}=\mathbb{R}^{n}$.
+As a result the tangent space $\mathcal{T}_{p}$, which is a linear approximation of the manifold at $p$, is simply $\mathbb{R}^{n}$ for all $p$.
+The metric tensor $g_{p}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is a smoothly varying positive-definite bilinear form on the tangent space.
 It defines an inner product in the tangent space, allowing us to measure lengths and angles.
-[…]
+For a tangent vector $v\in\mathcal{T}_{p}$, the norm is defined as $\|v\|=\sqrt{g_{p}(v,v)}$, providing a notion of length in the tangent space.
+Just as we measure the length of a curve in Euclidean space by integrating its speed, the length of a smooth curve $\gamma(t)$ on the manifold is given by $\int\|\dot{\gamma}(t)\|\,\mathrm{d}t$, where $\dot{\gamma}(t)$ is the derivative of $\gamma(t)$ with respect to $t$.
+This integral accumulates the infinitesimal lengths along the curve, accounting for how the metric varies across the manifold.
 The distance between two points on the manifold is defined as the minimal length of a curve connecting the points.
 The curves that minimize this length are known as geodesics, which are the generalization of straight lines in Euclidean space to curved spaces.
+These geodesics satisfy a second-order differential equation known as the geodesic equation (see Section A.1).
+In our context, these geodesics will represent optimal paths through the data space that respect the underlying probability density structure.
 2.2 Fermat Distances
-Density-based distances (DBDs) provide a way to define distances between points in a distribution by utilizing the underlying probability density function.
-This approach can be particularly useful for tasks such as clustering and pathfinding in data analysis and machine learning.
+Having established the framework of Riemannian geometry, we now turn to density-based distances (DBDs), which provide a way to define distances between points sampled from a distribution by utilizing the underlying probability density function.
+This approach can be particularly useful for tasks such as clustering and pathfinding.
 If designed well, a DBD can capture desirable properties such as:
@@
 In the language of Riemannian geometry, we define a metric tensor of the form
-$g(u,v)=\frac{\langle u,v\rangle}{f(p)^{2}},$ (1)
+$g_{x}(u,v)=\frac{\langle u,v\rangle}{f(p(x))^{2}},$ (1)
 […]
@@
 […]
-3.2.1 Fermat distances which scale with dimension
+This integral accumulates the Euclidean length element $\|\dot{\gamma}(t)\|$, weighted inversely by the density along the path.
+As a result, segments of the path that pass through low-density regions contribute more to the total length than those through high-density regions.
+Fermat Distances Which Scale with Dimension
 Figure 2: Geodesics of the standard normal distribution, projected into 2D, taken between two orthogonal points at distance $\sqrt{D}$ from the origin.
 The paths are scaled by $\sqrt{D}$ to enable comparison.
-(Left) Unscaled metric (same temperature for all dimensions) leads to sharply curved geodesics in high dimensions.
-(Right) Scaled metric (temperature equal to dimension) leads to consistent geodesics in all dimensions (note that the lines are overlapping).
-If $\beta=1$, this leads to unintuitive behavior in higher dimensions.
+Left: Unscaled metric (same temperature for all dimensions) leads to sharply curved geodesics in high dimensions.
+Right: Scaled metric (temperature equal to dimension) leads to consistent geodesics in all dimensions (note that the lines are overlapping).
+While the Fermat distance framework is theoretically elegant, setting $\beta=1$ (as is common in previous work) leads to unintuitive behavior in higher dimensions.
 Consider the standard normal distribution in $D$ dimensions.
@@
 Therefore, geodesics in the $D$-dimensional standard normal distribution are the same as in the two-dimensional standard normal distribution, once we account for the rotation into the plane of $x_{1}$ and $x_{2}$.
-However, despite this similarity, the typical distance from the origin of points sampled from the $D$-dimensional standard normal distribution is much larger, with a mean of $\sqrt{D}$.
-As a result, geodesics will start and end in very low-density regions relative to the two-dimensional standard normal.
-This means that most of the trajectory is spent moving towards and then outwards from the high-density region near the origin.
-To address this issue, if we scale the temperature linearly with dimension ($\beta=1/D$), the results become consistent with the two-dimensional case (see Fig.
-2).
-Scaling the temperature with dimension also increases numerical stability.
-This is because the effect of the probability density is reduced, leading to more Euclidean-like geodesics.
-This effect is clearly visible in Fig.
-2, where the geodesics are much closer to straight lines in the scaled case.
+However, despite this geometric similarity, there’s a crucial scaling issue: the typical distance from the origin of points sampled from the $D$-dimensional standard normal distribution is much larger, with a mean of $\sqrt{D}$.
+As a result, geodesics will start and end in exponentially low-density regions relative to the two-dimensional standard normal.
+This forces most of the trajectory to make a dramatic detour towards and then away from the high-density region near the origin, creating increasingly extreme curved paths as dimension grows.
+To address this issue, if we scale the temperature linearly with dimension ($\beta=1/D$), the results become consistent with the two-dimensional case (see Figure 2).
+In Section A.3, we explore this argument in more detail and generalize it to non-Gaussian distributions.
+This scaling has two key benefits: First, it ensures that geodesic behavior remains intuitive across dimensions, with similar qualitative properties regardless of the ambient dimension.
+Second, it significantly improves numerical stability by reducing the effect of the probability density, leading to more Euclidean-like geodesics.
+This effect is clearly visible in Figure 2, where the geodesics are much closer to straight lines in the scaled case.
 2.3 Geodesic Equation
-From the metric, we can compute the geodesic equation.
-This is the second-order differential equation which all paths of minimum length (geodesics) must satisfy.
-In our case, we have the relatively simple solution for a trajectory $\gamma(t)$:
+Figure 3: Comparison of geodesic parameterizations.
+Left: Constant metric speed leads to varying Euclidean step sizes, making numerical relaxation challenging.
+Right: Constant Euclidean speed enables uniform discretization and more stable numerics.
+Points along the paths indicate discretization steps.
+Note the slightly different solutions: the left-hand path is approximately 0.1% longer.
+From the metric, we can compute the geodesic equation, which characterizes the paths of minimum length between points.
+Just as a straight line in Euclidean space can be described by a second-order differential equation ($\ddot{x}=0$), geodesics in our curved space satisfy a more complex equation (see Theorem A.2).
+For a geodesic trajectory $\gamma(t)$, we have:
 $\ddot{\gamma}-2\beta(s(\gamma)\cdot\dot{\gamma})\dot{\gamma}+\beta s(\gamma)\|\dot{\gamma}\|^{2}=0$ (3)
@@
 The geodesic equation has the property that the speed is constant with respect to the metric, meaning $\sqrt{g(\dot{\gamma},\dot{\gamma})}$ is constant.
-In other words, the geodesic distance traveled in a given time interval is the same at all points of the curve.
-If $\gamma$ passes from a high-probability region to a low-probability one, the distances in the low-probability region could be orders of magnitude larger than in the high-probability region.
-This makes numerically solving the equations challenging due to the vastly different Euclidean velocities along the trajectory.
-It is convenient to reparameterize the equation such that the Euclidean velocity is constant instead, as this is much easier to handle numerically.
-This is especially useful for solving via relaxation, where having equal-length segments leads to much more stable behavior.
-Making this change results in a very similar equation (removing only the factor of 2 in the second term):
+While mathematically elegant, this creates numerical challenges: when $\gamma$ passes from a high-probability region to a low-probability one, the Euclidean velocities can differ by orders of magnitude.
+For example, if the probability density drops by a factor of 1000, the Euclidean velocity must increase by the same factor to maintain constant metric speed.
+This dramatic speed variation makes standard numerical integrators unstable and requires extremely small step sizes in low-density regions while potentially overshooting in high-density regions.
+To address these numerical challenges, we reparameterize the equation to maintain constant Euclidean speed instead of constant metric speed (see Theorem A.3).
+This simplification is particularly valuable for numerical methods like relaxation, where uniform spacing of points along the trajectory leads to better stability.
+Figure 3 illustrates how this reparameterization enables uniform discretization compared to the varying step sizes required by constant metric speed.
+The reparameterized equation differs only slightly from the original (losing just the factor of 2 in the second term):
 $\ddot{\varphi}-\beta(s(\varphi)\cdot\dot{\varphi})\dot{\varphi}+\beta s(\varphi)\|\dot{\varphi}\|^{2}=0$ (4)
-See Section A.1 for derivations of these results.
-4 Methods
-4.1 Ground truth distances through relaxation
-A relaxation scheme allows one to solve a differential equation constrained by boundary conditions by iteratively updating the solution such that it better obeys the differential equation.
-We divide the interval $[0,1]$ into equal time steps of size $h=1/n$ and write $\varphi_{i}=\varphi(ih)$.
-Initialize the values of the curve $\varphi$ at $t=ih$ for $i=0,\dots,n$, with the constraint that $\varphi_{0}$ and $\varphi_{n}$ are the chosen endpoints.
-The initial curve could be the straight line connecting the end points, or a more informative guess if additional information is available.
-Equation 4 leads to the following relaxation scheme. Define
-$v_{i}=\frac{\varphi_{i+1}-\varphi_{i-1}}{2}$ (5)
-then update with
-$\varphi^{\prime}_{i}=\frac{\varphi_{i+1}+\varphi_{i-1}}{2}+\frac{\beta}{2}\left(s(\varphi_{i})\|v_{i}\|^{2}-(s(\varphi_{i})\cdot v_{i})v_{i}\right)$ (6)
-See Section A.2 for a derivation.
-4.1.1 Implementation
-In practice, some additional features are needed to ensure stable convergence of the relaxation scheme, especially when the score function is not particularly smooth.
-Firstly, we found that oscillatory features commonly arise.
-To address this, we use smoothing after the update, returning a value which is interpolated between the value at the node and the average of its neighbors.
-Secondly, for rougher score functions we use a step size between 0 and 1, where 0 corresponds to no update and 1 corresponds to the ordinary relaxation step.
-See Algorithm 1.
-Algorithm 1 One step of relaxation.
-$\varphi=\{\varphi_{i}\}_{i=0}^{n}$ is the current state of the trajectory, $s$ is the score function, $\beta$ the inverse temperature, $\gamma\in[0,1)$ the smoothing parameter and $\alpha\in(0,1]$ the step size.
-This step should be iterated until convergence, which can be measured via the difference between $\varphi$ and $\varphi^{\prime}$.
-function RelaxationStep($\varphi,s,\beta,\gamma,\alpha$)
-     $v_{i}\leftarrow\frac{1}{2}(\varphi_{i+1}-\varphi_{i-1})$triangleright $i=1,\dots,n-1$
-     $\varphi^{\prime}_{i}\leftarrow\frac{1}{2}(\varphi_{i+1}+\varphi_{i-1})+\frac{\beta}{2}\left(s(\varphi_{i})\|v_{i}\|^{2}-(s(\varphi_{i})\cdot v_{i})v_{i}\right)$
-     $\varphi^{\prime}_{i}\leftarrow(1-\gamma)\varphi^{\prime}_{i}+\frac{\gamma}{2}(\varphi^{\prime}_{i+1}+\varphi^{\prime}_{i-1})$
-     $\varphi^{\prime}_{i}\leftarrow\alpha\varphi^{\prime}_{i}+(1-\alpha)\varphi_{i}$
-     return $\varphi^{\prime}$
-end function
+We stress that, while Equation 3 is easily obtained from standard results on conformal metrics (e.g., Carroll (2004, Appendix G)), the reparameterized form in Equation 4 is, to our knowledge, a novel contribution.
 2.4 Weighted Graph Algorithm Based on Euclidean Distance
+Figure 4: Illustration of the power-weighted graph approximation.
+Left: Sample points (dots) and density (gradient) from a 2D standard normal.
+Center: k-nearest neighbor connections between points, with edge weights determined by cubic Euclidean distances.
+Right: A shortest path (solid blue) found by Dijkstra’s algorithm compared to the true geodesic (dashed red).
+Note how the discrete path approximates the continuous geodesic by following high-density regions, but is not very smooth.
+A practical approach to computing Fermat distances is to discretize the problem using sample points.
 If we have a large number of samples from $p(x)$, we can approximate the density between close neighbors $x_{1}$ and $x_{2}$ as constant and hence the shortest path between them as a straight line.
-The density is roughly proportional to the inverse of the Euclidean distance, raised to the power $d$, where $d$ is the intrinsic dimension of the distribution (Bijral et al., 2012):
+Following Bijral et al.
+(2011), the density is roughly proportional to the inverse of the Euclidean distance, raised to the power $d$, where $d$ is the intrinsic dimension of the distribution:
 $p(\gamma)\propto\frac{1}{\|x_{1}-x_{2}\|^{d}}$ (5)
@@
 $\operatorname{dist}(x_{1},x_{2})\propto\|x_{1}-x_{2}\|^{\beta d+1}$ (6)
-Bijral et al.
-(2012) therefore proposes to approximate shortest paths between more distant points by constructing a $k$-nearest neighbors graph of the samples with edge weights given by $\|x_{1}-x_{2}\|^{\beta d+1}$ and applying Dijkstra’s algorithm.
+This elegant relationship suggests a simple algorithmic approach: Bijral et al.
+(2011) proposes to approximate shortest paths between more distant points by constructing a $k$-nearest neighbors graph of the samples with edge weights given by $\|x_{1}-x_{2}\|^{\beta d+1}$ and applying Dijkstra’s algorithm.
+Figure 4 illustrates the method.
 Hwang et al.
@@
 While this approach is appealing in its simplicity, can be easily implemented using standard libraries, and comes with theoretical guarantees, the density estimate is too poor to provide accurate estimates of Fermat distances in practice.
-In the experimental section 5.1.1 we show that paths computed with this approach do not converge to the ground truth geodesics as sample size increases, even in very simple 2-dimensional distributions such as the standard normal.
-We are unsure why we observe such a large deviation from the theoretical prediction.
-It is possible that the unknown constant $\theta_{0}$ is very small, leading to extremely slow convergence, despite the exponential function.
-We leave investigation into this point to future work.
+In the experimental section 4.1.1 we show that paths computed with this approach converge extremely slowly to the ground truth geodesics as sample size increases, even in very simple 2-dimensional distributions such as the standard normal.
+This significant gap between theoretical guarantees and practical performance leads us to speculate that the unknown constant $\theta_{0}$ is very small, resulting in extremely slow convergence despite the exponential form of the bound.
+This observation motivates our development of more accurate density estimation methods in the following sections.
+We leave a detailed theoretical investigation of this convergence behavior to future work.
 2.5 Normalizing Flows
-Normalizing flows (Rezende and Mohamed, 2015; Kobyzev et al., 2020; Papamakarios et al., 2021) are a set of techniques to learn a probability density via a parameterized diffeomorphism $f_{\theta}$ which maps from data to latent space.
+Normalizing flows (Rezende & Mohamed, 2015; Kobyzev et al., 2020; Papamakarios et al., 2021) provide a powerful framework for learning complex probability distributions.
+The key idea is to transform a simple base distribution (like a standard normal) into the target distribution through a series of invertible transformations.
+Formally, this is achieved via a parameterized diffeomorphism $f_{\theta}$ which maps from data to latent space.
 The probability density is obtained through the change of variables formula:
@@
 where $\pi$ is a simple latent distribution.
-A common strategy to design normalizing flows is through coupling blocks (Dinh et al., 2014, 2016; Durkan et al., 2019), where $f_{\theta}$ is the composition of a series of blocks, each with a triangular Jacobian.
+A common strategy to design normalizing flows is through coupling blocks (Dinh et al., 2014, 2017; Durkan et al., 2019), where $f_{\theta}$ is the composition of a series of blocks, each with a triangular Jacobian.
 This makes the determinant easy to calculate, and hence the above formula is tractable.
@@
 $\mathcal{L}=\mathbb{E}_{p_{\text{data}}(x)}[-\log p_{\theta}(x)]$ (8)
-4.4 Density-weighted graph algorithm
-In order to better estimate probability densities, we use a normalizing flow trained on the data and directly use the density predicted by the model.
-In order to approximate the geodesic distance between close neighbors, we assume that the geodesic connecting them is approximately a straight line, and estimate its length by a discrete approximation:
-$\operatorname{dist}(x_{1},x_{2})\approx\sum_{i=1}^{S}\frac{\|y_{i}-y_{i-1}\|}{p_{\theta}(y_{i-1/2})}$ (11)
-Here the $y_{i}$ are points which interpolate $x_{1}$ and $x_{2}$ with $y_{0}=x_{1}$ and $y_{S}=x_{2}$, and $y_{i-1/2}$ is the midpoint of $y_{i-1}$ and $y_{i}$.
-This allows us to take a k-nearest neighbors graph of the data and replace the edge weights by the approximate distance.
-This results in a sparse graph, to which we can apply Dijkstra’s algorithm to estimate shortest paths and distances between any two points.
-See Algorithm 2.
-Algorithm 2 Construction of a density-weighted graph which can subsequently be used to compute shortest paths and distances between points.
-$X$ is the data matrix, $p$ the probability density, $\beta$ the inverse temperature, $k$ the number of neighbors, and $S$ the number of segments with which to approximate the integral.
-The computations can also be performed in the log space, returning log edge weights, which is more numerically stable in higher dimensions.
-function DensityGraph($X,p,\beta,k,S$)
-     $\mathcal{G}\leftarrow\texttt{KNN}(X,k)$triangleright Construct $k$-nearest neighbor graph
-     for $(l,m)\in\mathcal{G}$ $\triangleright$ Iterate over edges
-         $y_{i}\leftarrow(1-\frac{i}{S})X_{l}+\frac{i}{S}X_{m}$$\triangleright$ For $i=0,\dots,S$
-         $y_{i-1/2}\leftarrow\frac{1}{2}(y_{i-1}+y_{i})$$\triangleright$ For $i=1,\dots,S$
-         $\mathcal{G}_{lm}\leftarrow\sum_{i=1}^{S}\|y_{i}-y_{i-1}\|/p(y_{i-1/2})^{\beta}$
-     end for
-     return $\mathcal{G}$
-end function
-4.5 Relaxation with a learned score model
-The score function is the gradient of the log-likelihood with respect to spatial inputs, i.e., $s(x)=\nabla_{x}\log p(x)$.
-Note the difference from the definition commonly used in statistic, where the derivative is with respect to model parameters.
-Score matching (Hyvärinen, 2005) is a family of methods to estimate this function from data, including variants such as sliced score matching (Song et al., 2020a), denoising score matching (Vincent, 2011) and diffusion models (Song et al., 2020b).
+2.6 Score Matching
+While normalizing flows provide a way to estimate the density directly, our geodesic equations actually depend on the score function $s(x)=\nabla_{x}\log p(x)$, which is the gradient of the log-likelihood with respect to spatial inputs.
+Although one could obtain scores by differentiating the density model, we find that these derived scores are too noisy to be used effectively in practice.
+We speculate this reflects an inherent trade-off between optimal learning of the log density and its gradient, which we illustrate with simple examples in Section C.3.
+Therefore, we turn to score matching (Hyvärinen, 2005), which offers a direct way to estimate the score function from data, bypassing the need to estimate the density itself.
+This family of methods includes several variants such as sliced score matching (Song et al., 2020), denoising score matching (Vincent, 2011) and diffusion models (Song et al., 2021).
+For our purposes, we use sliced score matching due to its simplicity and ability to estimate the true score, rather than a noisy score as in denoising score matching or diffusion methods.
 Sliced Score Matching
-We can learn the score of $p(x)$ by minimizing the following objective (known as the score matching objective):
+We can learn the score of $p(x)$ by minimizing the score matching objective:
 $\mathbb{E}_{p(x)}[\operatorname{tr}(\nabla_{x}s_{\theta}(x))+\tfrac{1}{2}\|s_{\theta}(x)\|^{2}]$ (9)
@@
 where $v$ is sampled from an appropriate distribution, typically standard normal.
-Since only the score is needed for the relaxation algorithm described in Section 4.1, we can use a trained score model to perform relaxation to approximate geodesics.
-In practice, we initialize the path from a graph-based method to speed up convergence and avoid local minima.
+3 Methods
+Figure 5: Illustration of geodesic approximations by graph-based methods (left) and relaxation methods (right) on 200 points sampled from the 2D standard normal.
+Left: the power-weighted shortest path method (Section 2.4) produces the roughest path.
+The shortest path from an inverse-density weighted graph using a model density (Section 3.2) is more direct, but not quite as close to the true geodesic as the equivalent method using the ground truth density.
+The model is defined as Gaussian with the same mean and covariance as the data.
+Right: relaxation (Algorithm 1) on the model score function results in a very close approximation to the true geodesic.
+3.1 Ground Truth Distances through Relaxation
+We can solve Equation 4 in several ways to find geodesics between two points.
+The shooting method starts at one point with an initial velocity guess and integrates the equation forward, iteratively adjusting the initial velocity to reach the target point.
+Alternatively, the relaxation method discretizes the interval $[0,1]$ into $n$ equal time steps of size $h=1/n$, writing $\varphi_{i}=\varphi(ih)$ for the curve values at these points.
+Starting from an initial path (either a straight line between endpoints or a more informed guess), it gradually updates the intermediate points while keeping the endpoints fixed to better satisfy the differential equation.
+Each method has its advantages: shooting is more efficient for exploring geodesics from a starting point, while relaxation is more robust for finding paths between specific endpoints, especially in complex density landscapes where shooting might miss the target.
+We opt to use relaxation.
+Equation 4 leads to the relaxation scheme given in Algorithm 1, which updates each point along the curve based on its neighbors and the local score function.
+The random ordering of updates helps avoid oscillatory solutions and improves convergence.
+See Appendix A for a detailed derivation.
+3.2 Density-Weighted Graph Algorithm
+To improve upon previous graph-based methods, we replace their density estimates with those from a normalizing flow trained on the data.
+For nearby points, we can approximate the geodesic connecting them as a straight line and estimate its length through numerical integration.
+Specifically, given two points $x_{1}$ and $x_{2}$, we estimate their distance as:
+$\operatorname{dist}(x_{1},x_{2})$ $\approx\sum_{i=1}^{S}\frac{\|y_{i}-y_{i-1}\|}{p_{\theta}(y_{i-1/2})^{\beta}}$ (11)
+$=\frac{\|x_{1}-x_{2}\|}{S}\sum_{i=1}^{S}\frac{1}{p_{\theta}(y_{i-1/2})^{\beta}}$ (12)
+where $\{y_{i}\}_{i=0}^{S}$ are equally spaced points along the line from $x_{1}$ to $x_{2}$ (i.e., $y_{0}=x_{1}$ and $y_{S}=x_{2}$), and $y_{i-1/2}$ denotes the midpoint between consecutive points.
+We use this distance measure to construct a k-nearest neighbors graph, with edge weights given by the approximate geodesic distances rather than Euclidean distances (see Algorithm 2).
+The resulting sparse graph structure allows us to efficiently compute distances between any pair of points using Dijkstra’s shortest path algorithm.
+Due to the large dynamic range of the density values involved, we perform all distance calculations and graph operations in log space to maintain numerical stability.
+Algorithm 1 One step of relaxation (iterate until convergence)
+  Input: trajectory $\varphi=\{\varphi_{i}\}_{i=0}^{n}$, score function $s$, inverse temperature $\beta$
+  $I=\{1,\dots,n-1\}$
+  $\mathtt{shuffle}(I)$
+  for $i$ in $I$ do
+     $v_{i}=\frac{1}{2}(\varphi_{i+1}-\varphi_{i-1})$
+     $w_{i}=\frac{1}{2}\beta[s(\varphi_{i})\|v_{i}\|^{2}-(s(\varphi_{i})\cdot v_{i})v_{i}]$
+     $\varphi_{i}=\frac{1}{2}(\varphi_{i+1}+\varphi_{i-1})+w_{i}$
+  end for
+  Return: $\varphi$
+Algorithm 2 Construction of density-weighted graph
+  Input: data matrix $X$, density function $p$, inverse temperature $\beta$, neighbors $k$, segments $S$
+  $\mathcal{G}=\mathtt{KNN}(X,k)$Construct k-nearest neighbor graph}
+  for edge $(l,m)$ in $\mathcal{G}$ do
+     $\Delta x=(X_{m}-X_{l})/S$
+     for $i=1$ to $S$ do
+        $y_{i-1/2}=X_{l}+(i-0.5)\Delta x$
+     end for
+     $\mathcal{G}_{lm}=\Delta x\sum_{i=1}^{S}1/p(y_{i-1/2})^{\beta}$
+  end for
+  Return: $\mathcal{G}$
+Figure 6: The length of paths in learned and ground-truth density weighted graphs (orange and blue) converge to the ground truth at roughly the same rate with increasing sample size, while the power-weighted graph method (green) converges at a far inferior rate.
+3.3 Relaxation with a Learned Score Model
+The relaxation algorithm described in Section 3.1 requires only the score function $s(x)$, not the density itself.
+This observation allows us to use a score model trained directly via score matching, bypassing potential issues with density estimation.
+In practice, we find this approach more effective than computing scores by differentiating a normalizing flow (see experimental results in Section 4).
+To ensure robust convergence, we initialize the path using the graph-based method from the previous section.
+This initialization provides a reasonable approximation of the geodesic, helping the relaxation process avoid local minima and converge more quickly to the optimal path.
 4 Experiments
+We have published the code to reproduce these experiments in the following GitHub repository: https://github.com/vislearn/Fermat-Distance.
 Performance Metric
-After solving for a shortest path $\varphi$ between two points $x_{1}$ and $x_{2}$, we compute its length $L(\varphi)$ by using the ground truth density function, then compare this value to the ground truth distance between the two points:
-$M(\varphi)=\log\frac{L(\varphi)}{\operatorname{dist}(x_{1},x_{2})}$ (14)
-Provided we have access to the ground truth density, this performance metric can be calculated independently of how $\varphi$ is obtained, and has a value of zero only if $\varphi$ is the ground truth geodesic.
-The metric will otherwise take positive values.
-5.1 Scaling the dataset
-Figure 3: The length of shortest paths in power-weighted distance graphs do not converge to the ground truth with increasing sample size (green).
-In contrast, learned and ground-truth density weighted graphs (orange and green) do converge, at similar rates.
-In this section we perform experiments on a set of 5 two-dimensional datasets for which we know the ground truth density (see Fig.
-3).
-Details on how the datasets were created can be found in Appendix B.
-5.1.1 Power-weighted graph paths do not converge with increasing sample size
-We first estimate paths using the power-weighted graph method (Bijral et al., 2012), applied to a k-nearest neighbor graph.
-We find that in all 5 datasets, increasing the sample size leads to saturating performance (see Fig.
-3, green lines).
-In Fig.
-6 in the appendix we show that increasing the number of neighbors does not improve performance.
-5.1.2 Poor convergence is due to poor density estimation
-In order to investigate the root cause of the poor convergence, we designed other graph methods which use the nearest neighbor estimator for the density, but do not use power-weighted graph edges.
-We use 4 variants for the density estimate (see Section C.1).
-We find that all 4 have very similar performance to the power-weighted graph, whereas the equivalent estimators using the ground truth density quickly converge.
-See Section C.1 for plots supporting this claim.
-5.1.3 Learning the density leads to convergent paths
-Since the problem lies in poor density estimation, we hope to fix the problem by better density estimation.
-To this end, we use a normalizing flow to learn the density from samples, then use Algorithm 2 to construct a graph over which we compute shortest paths.
-We find that the paths obtained converge at close to the same rate as paths obtained by an equivalent method using the ground truth densities (Fig.
-3 orange and blue lines).
-Full experimental details can be found in Appendix B.
+To evaluate the quality of computed paths, we introduce the Log Path Ratio (LPR) metric that can be applied to any path $\varphi$ connecting two points $x_{1}$ and $x_{2}$.
+We compute the path length $L(\varphi)$ using the ground truth density and compare it to the true geodesic distance:
+$\operatorname{LPR}(\varphi)=\log\frac{L(\varphi)}{\operatorname{dist}(x_{1},x_{2})}$ (13)
+This metric is zero only for the true geodesic and positive otherwise, providing a principled way to evaluate different methods when ground truth densities are available.
+4.1 Convergence Analysis on 2D Datasets
+To evaluate convergence behavior, we analyze five two-dimensional datasets with known ground truth densities, ranging from simple unimodal distributions to complex multimodal mixtures (shown in Figure 6).
+For each dataset, we compute the average LPR over 1000 randomly sampled paths.
+Full experimental details and dataset descriptions are provided in the supplementary material.
+4.1.1 Poor Convergence of Power-Weighted Methods
+We first examine the power-weighted graph method of Bijral et al.
+(2011), which constructs paths on k-nearest neighbor graphs.
+Despite theoretical guarantees, we find that performance improves extremely slowly with increasing sample size across all datasets (Figure 6, green lines).
+4.1.2 Density Estimation is the Bottleneck
+To identify the source of this poor convergence, we developed several variant methods that use different nearest-neighbor density estimators while maintaining the same graph structure (see supplementary material for details).
+All variants show similarly poor performance, while their counterparts using ground truth densities converge rapidly.
+This clearly identifies density estimation as the primary bottleneck.
+4.1.3 Improved Convergence through Better Density Estimation
+Having identified poor density estimation as the core issue, we replace nearest-neighbor estimates with a learned normalizing flow model.
+Using this density estimate in Algorithm 2, we find that the resulting paths converge at nearly the same rate as those computed using ground truth densities (Figure 6, orange and blue lines).
+This demonstrates that modern density estimation techniques can effectively close the gap between theoretical guarantees and practical performance.
+Full experimental details can be found in the supplementary material.
 4.2 Scaling the Dimension
-Figure 4: Mean log error with a sample size of 200,000.
-Graph-based paths (the first three methods) converge very slowly in higher dimensions with increasingly worse performance as dimension increases.
-When using the scaled metric ($\beta=1/D$), convergence improves over the unscaled metric, but all graph-based methods perform similarly in higher dimensions.
-Doing relaxation with a learned score function (red line) leads to much faster convergence, avoiding the curse of dimensionality.
-Machine learning datasets often lie in high-dimensional spaces, so it is crucial that any method applied to them scales well, avoiding the curse of dimensionality as much as possible.
-We test this by taking one of our two-dimensional datasets, namely the standard normal, and scaling it up to 25 dimensions.
-We perform experiments both for the scaled and unscaled version of the metric.
-5.2.1 Graph-based paths do not converge well in higher dimensions
-Our first finding is that all graph-based methods show declining performance as the dimension increases, for both unscaled and scaled metrics (Fig.
-4).
-This can be straightforwardly explained by the curse of dimensionality: with increasing dimension but fixed sample size, it is increasingly unlikely that points in the dataset lie close to the true trajectory.
-The shortest path in the graph becomes increasingly rough in contrast to the smooth ground truth trajectory.
-5.2.2 Relaxation using a learned score function improves convergence rates
-In order to obtain smoother paths, we turn to the same relaxation algorithm we use to solve for ground truth geodesics.
-Instead of using the ground truth score function, we use a learned score function.
-An obvious choice would be to use automatic differentiation to get the score from a trained normalizing flow, but we find the resulting scores are too noisy to be useful, inhibiting reliable convergence of the relaxation scheme.
-Instead, we learn a score model with sliced score matching, which results in a smoother function suitable for performing relaxation.
-Full experimental details can be found in Appendix B.
-We find that this fixes the convergence problems, both for the unscaled and scaled metrics, though convergence is more reliable in the scaled case (Fig.
-4).
-Please see Fig.
-8 and Fig.
-9 in the appendix for more detailed evidence.
+Figure 7: Mean LPR with a sample size of 200,000.
+Graph-based methods (first three) show rapidly degrading performance as dimension increases, even with the scaled metric ($\beta=1/D$).
+In contrast, relaxation with a learned score function (red line) maintains good performance across dimensions, effectively avoiding the curse of dimensionality.
+Machine learning datasets often lie in high-dimensional spaces, making dimensional scaling behavior crucial for practical applications.
+We investigate this by extending one of our two-dimensional datasets (the standard normal) up to 25 dimensions, testing both the scaled ($\beta=1/D$) and unscaled ($\beta=1$) versions of the metric.
+4.2.1 Graph-Based Methods Fail in Higher Dimensions
+All graph-based methods show rapidly degrading performance as dimension increases, regardless of whether we use scaled or unscaled metrics (Figure 7).
+This failure is a direct consequence of the curse of dimensionality: with fixed sample size but increasing dimension, the probability of finding sample points near the true geodesic becomes vanishingly small.
+As a result, graph-based paths become increasingly jagged approximations of the smooth ground truth trajectory.
+4.2.2 Score-Based Relaxation Overcomes Dimensional Scaling
+To address this fundamental limitation of graph-based methods, we apply our relaxation algorithm (Algorithm 1) with a learned score function.
+While one might expect to obtain scores by differentiating a trained normalizing flow, we find these derived scores too noisy for reliable convergence (see Section C.3 for analysis of this phenomenon).
+Instead, direct score estimation via sliced score matching provides smoother gradients that enable stable relaxation.
+This score-based approach dramatically improves performance, maintaining good convergence across dimensions for both scaled and unscaled metrics, though the scaled version ($\beta=1/D$) shows more reliable behavior (Figure 7).
+Detailed experimental results and analysis can be found in the supplementary material.
+4.3 MNIST
+We also perform preliminary experiments on MNIST within the latent space of an autoencoder, as described in Section C.1.
+We find that we obtain interpretable and reasonable geodesics connecting digits within this space, and that digits cluster together in an unsupervised way.
 5 Conclusion
-In conclusion, our study focuses on understanding the convergence properties of various algorithms to ground-truth geodesics by employing toy distributions with known densities.
-We investigate the standard normal distribution in dimensions higher than 2, highlighting the challenges faced by graph-based methods even in this simple scenario.
-Our findings demonstrate significant improvements through the use of normalizing flows and score models, yet further research is needed to address more complex distributions and real-world data where the underlying density is unknown.
-Future work should aim to unify the roles of normalizing flows and score models into a single cohesive model, enhancing the method’s applicability and efficiency.
-Additionally, a deeper theoretical analysis is required to understand why our proposed methods converge effectively, while previous methods with consistency guarantees do not.
-This work lays the groundwork for practical applications of density-based distances, setting the stage for more robust and scalable solutions in metric learning.
+We have presented a systematic study of geodesic computation in density-based metrics, using distributions with known ground truth to rigorously evaluate different approaches.
+Our investigation reveals several key insights:
+First, we demonstrate that previous graph-based methods, despite their theoretical guarantees, converge impractically slowly even for simple distributions like the standard normal.
+Second, we show that this poor convergence stems primarily from inadequate density estimation rather than the graph approximation itself.
+Third, we establish that modern deep learning approaches—normalizing flows for density estimation and score matching for gradients—can dramatically improve performance.
+However, significant challenges remain.
+While our methods work well on toy distributions with known densities, extending them to complex real-world data requires further research.
+Key directions for future work include:
+  • Unifying normalizing flows and score models into a single framework to improve efficiency
+  • Developing theoretical understanding of why our methods converge effectively while previous approaches with consistency guarantees do not
+  • Extending these methods to higher-dimensional densities typical in real applications
+This work establishes a foundation for practical density-based distances, bringing theoretical elegance closer to practical applicability in metric learning.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Acknowledgments
+This work is supported by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2181/1 - 390900948 (the Heidelberg STRUCTURES Cluster of Excellence).
+The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG.
 Appendix A Derivations
 A.1 Geodesic Equation
-The geodesic equation which characterizes paths of zero acceleration $\gamma$ is
-$\ddot{\gamma}^{i}+\Gamma^{i}_{jk}\dot{\gamma}^{j}\dot{\gamma}^{k}=0$ (15)
-where $\Gamma^{i}_{jk}$ are the Christoffel symbols of the second kind
+We begin by deriving the geodesic equation for our conformal metric.
+This is a well-known result in differential geometry (see e.g., Carroll (2004, Appendix G)).
+The key steps are computing the Christoffel symbols and then simplifying the resulting equation.
+Lemma A.1 (Christoffel Symbols for Conformal Metrics).
+For a conformal metric of the form $g_{ij}=\lambda^{2}\delta_{ij}$, the Christoffel symbols are:
+$\Gamma^{i}_{jk}=\frac{\partial\log\lambda}{\partial x^{k}}\delta^{i}_{j}+\frac{\partial\log\lambda}{\partial x^{j}}\delta^{i}_{k}-\frac{\partial\log\lambda}{\partial x^{l}}\delta^{il}\delta_{jk}$ (14)
+Proof.
+The Christoffel symbols are given by:
 $\Gamma^{i}_{jk}=\frac{1}{2}g^{il}\left(\frac{\partial g_{lj}}{\partial x^{k}}+\frac{\partial g_{lk}}{\partial x^{j}}-\frac{\partial g_{jk}}{\partial x^{l}}\right)$ (15)
-and $g^{il}$ is the inverse of $g$, i.e. such that $g^{il}g_{lj}=\delta^{i}_{j}$.
-Our metric has the form $g_{ij}=\lambda^{2}\delta_{ij}$, meaning that
+For our metric:
 $\Gamma^{i}_{jk}$ $=\frac{1}{2\lambda^{2}}\delta^{il}\left(\frac{\partial\lambda^{2}}{\partial x^{k}}\delta_{lj}+\frac{\partial\lambda^{2}}{\partial x^{j}}\delta_{lk}-\frac{\partial\lambda^{2}}{\partial x^{l}}\delta_{jk}\right)$ (16)
 $=\delta^{il}\left(\frac{\partial\log\lambda}{\partial x^{k}}\delta_{lj}+\frac{\partial\log\lambda}{\partial x^{j}}\delta_{lk}-\frac{\partial\log\lambda}{\partial x^{l}}\delta_{jk}\right)$ (17)
-$=\frac{\partial\log\lambda}{\partial x^{k}}\delta^{i}_{j}+\frac{\partial\log\lambda}{\partial x^{j}}\delta^{i}_{k}-\frac{\partial\log\lambda}{\partial x^{l}}\delta^{il}\delta_{jk}$ (19)
-In this case $\lambda=p^{-\beta}$ and hence $\frac{\partial\log\lambda}{\partial x}=-\beta\frac{\partial\log p}{\partial x}=-\beta s$, where $s=\frac{\partial\log p}{\partial x}$ is the score function of the probability density.
-As a result
+The result follows by simplifying the Kronecker delta products. ∎
+Theorem A.2 (Geodesic Equation).
+For a metric of the form $g_{ij}=p^{-2\beta}\delta_{ij}$, where $p$ is a probability density and $\beta$ is a constant, the geodesic equation in vector notation is:
+$\ddot{\gamma}-2\beta(s(\gamma)\cdot\dot{\gamma})\dot{\gamma}+\beta s(\gamma)\|\dot{\gamma}\|^{2}=0$ (18)
+where $s=\nabla\log p$ is the score function of the probability density.
+Proof.
+For our metric, $\lambda=p^{-\beta}$ and hence $\frac{\partial\log\lambda}{\partial x}=-\beta\frac{\partial\log p}{\partial x}=-\beta s$. Substituting into the formula for Christoffel symbols:
 $\Gamma^{i}_{jk}=-\beta s_{k}\delta^{i}_{j}-\beta s_{j}\delta^{i}_{k}+\beta s_{i}\delta^{il}\delta_{jk}$ (19)
-Substituting into the geodesic equation leads to
+The geodesic equation $\ddot{\gamma}^{i}+\Gamma^{i}_{jk}\dot{\gamma}^{j}\dot{\gamma}^{k}=0$ then becomes:
 $\ddot{\gamma}^{i}-\beta s_{k}\dot{\gamma}^{i}\dot{\gamma}^{k}-\beta s_{j}\dot{\gamma}^{j}\dot{\gamma}^{i}+\beta s_{l}\delta^{il}\delta_{jk}\dot{\gamma}^{j}\dot{\gamma}^{j}=0$ (20)
-which can be written in vector notation as
-$\ddot{\gamma}-2\beta(s(\gamma)\cdot\dot{\gamma})\dot{\gamma}+\beta s(\gamma)\|\dot{\gamma}\|^{2}=0$ (22)
-Let $f:[0,1]\rightarrow[0,1]$ be a strictly increasing diffeomorphism.
-Given a curve $\gamma:[0,1]\rightarrow\mathbb{R}^{n}$, consider the curve $\varphi:[0,1]\rightarrow\mathbb{R}^{n}$ defined such that $\varphi(u)=\gamma(f^{-1}(u))$ or equivalently $\varphi(f(t))=\gamma(t)$.
-The derivatives of the two curves are related by
+which simplifies to the desired vector form. ∎
+Theorem A.3 (Constant Euclidean Speed Parameterization).
+Under a reparameterization that maintains constant Euclidean speed, the geodesic equation becomes:
+$\ddot{\varphi}-\beta(s(\varphi)\cdot\dot{\varphi})\dot{\varphi}+\beta s(\varphi)\|\dot{\varphi}\|^{2}=0$ (21)
+Proof.
+Let $f:[0,1]\rightarrow[0,1]$ be a strictly increasing diffeomorphism and define $\varphi(u)=\gamma(f^{-1}(u))$. The derivatives are related by:
 $\dot{\varphi}\dot{f}$ $=\dot{\gamma}$ (22)
-and
-[…]
-where derivatives of $\varphi$ are with respect to $u=f(t)$. By substituting these expressions into the geodesic equation, we arrive at the equivalent equation for $\varphi$:
-$\ddot{\varphi}\dot{f}^{2}+\dot{\varphi}\ddot{f}-2\beta(s(\gamma)\cdot\dot{\varphi}\dot{f})\dot{\varphi}\dot{f}+\beta s(\gamma)\|\dot{\varphi}\dot{f}\|^{2}=0$ (25)
-We would like $\varphi$ to have a constant Euclidean speed (the magnitude but not necessarily the direction of the velocity is constant), meaning
+$\ddot{\varphi}\dot{f}^{2}+\dot{\varphi}\ddot{f}$ $=\ddot{\gamma}$ (23)
+For constant Euclidean speed, we require:
 $\frac{d}{du}\left(\frac{1}{2}\|\dot{\varphi}\|^{2}\right)=\dot{\varphi}\cdot\ddot{\varphi}=0$ (24)
-Multiplying the geodesic equation by $\dot{\varphi}$, this is equivalent to requiring
-$\|\dot{\varphi}\|^{2}\ddot{f}-2\beta(s(\gamma)\cdot\dot{\varphi})\|\dot{\varphi}\|^{2}\dot{f}^{2}+\beta(s(\gamma)\cdot\dot{\varphi})\|\dot{\varphi}\|^{2}\dot{f}^{2}=0$ (27)
-implying that
+Substituting into the geodesic equation and solving for $\ddot{f}$ yields:
 $\ddot{f}=\beta(s(\gamma)\cdot\dot{\varphi})\dot{f}^{2}$ (25)
-Resubstituting into the geodesic equation and dividing by $\dot{f}^{2}$ leads to the result:
-$\ddot{\varphi}-\beta(s(\varphi)\cdot\dot{\varphi})\dot{\varphi}+\beta s(\varphi)\|\dot{\varphi}\|^{2}=0$ (29)
+The result follows by substituting back and simplifying. ∎
 A.2 Relaxation Scheme
@@
 $\ddot{\varphi}_{i}\approx\frac{\varphi_{i+1}-2\varphi_{i}+\varphi_{i-1}}{h^{2}}$ (27)
-By substituting this into the differential equation we have
+By substituting this into the differential equation (Equation 21) we have
 $\frac{\varphi_{i+1}-2\varphi_{i}+\varphi_{i-1}}{h^{2}}-\beta(s(\varphi_{i})\cdot\dot{\varphi}_{i})\dot{\varphi}_{i}+\beta s(\varphi_{i})\|\dot{\varphi}_{i}\|^{2}\approx 0$ (28)
@@
 $\varphi_{i}=\frac{\varphi_{i+1}+\varphi_{i-1}}{2}+\frac{\beta}{2}\left(s(\varphi_{i})\|v_{i}\|^{2}-(s(\varphi_{i})\cdot v_{i})v_{i}\right)$ (31)
+A.3 Justification of Scaling Fermat Distances with Dimension
+We propose to scale the metric using a temperature equal to the dimension of the data.
+In order to justify this, consider the behavior of a generalized Gaussian distribution of the form $p(x)\propto\exp(-\|x\|^{\alpha}/\alpha)$.
+It can be shown that $\mathbb{E}[\|x\|^{\alpha}]=D$ and thus the typical distance from the origin of points sampled from $p$ scales as $D^{1/\alpha}$.
+The score in this case is $s(x)=-\|x\|^{\alpha-2}x$.
+Due to the rotational symmetry of the distribution, we can suppose without loss of generality that a sample $x$ lies in the plane of the first two coordinates and hence the problem reduces to the 2-dimensional one.
+However, since the typical distance of sampled points scales with the dimension, in order to avoid very exaggerated geodesics in high dimensions which deviate sharply towards the origin, we need to scale the density appropriately.
+By rescaling $\gamma$ in Equation 3 by $D^{1/\alpha}$, the score scales like $D^{(\alpha-1)/\alpha}$, and so the terms in the equation scale as:
+$\ddot{\gamma}-2\beta(s\cdot\dot{\gamma})\dot{\gamma}+\beta s\|\dot{\gamma}\|^{2}$ $\to D^{1/\alpha}\ddot{\gamma}-2\beta D^{(\alpha+1)/\alpha}(s\cdot\dot{\gamma})\dot{\gamma}+\beta D^{(\alpha+1)/\alpha}s\|\dot{\gamma}\|^{2}$ (32)
+$=D^{1/\alpha}\bigl{(}\ddot{\gamma}-2\beta D(s\cdot\dot{\gamma})\dot{\gamma}+\beta Ds\|\dot{\gamma}\|^{2}\bigr{)}$ (33)
+Therefore the geodesic equation can scale consistently if, and only if, $\beta=1/D$. The same reasoning applies to the reparameterized Equation 4.
+We can apply a similar argument to the Student-t distribution of the form $p(x)\propto(1+\|x\|^{2}/\nu)^{-(\nu+D)/2}$ with $\nu>2$.
+In the limit $\nu\to\infty$, the distribution becomes normal, and otherwise is characterized by heavier tails.
+The expected squared norm $\mathbb{E}[\|x\|^{2}]=\nu D/(\nu-2)$, so the typical distance from the origin scales like $\sqrt{D}$, as in a Gaussian distribution.
+The score function has the form $-\tfrac{\nu+D}{\nu+\|x\|^{2}}x$, and with $x\sim\sqrt{D}$, we have $s\sim x\sim\sqrt{D}$, again the same behavior as the Gaussian.
+Hence the scaling behavior is approximately the same as the Gaussian distribution, so the above argument for $\beta=1/D$ applies as well.
+Given that the same scaling behavior holds as in the generalized Gaussian, it suggests that $\beta=1/D$ is a natural default choice across a wide range of distributions.
 Appendix B Experimental Details
-Figure 5: The 5 two-dimensional datasets.
-Software libraries
-We used the numpy [Harris et al., 2020] and scipy [Virtanen et al., 2020] Python libraries for basic numerical and graph operations and matplotlib [Hunter, 2007] for plotting.
-We used scikit-learn [Pedregosa et al., 2011] to compute nearest neighbor graphs and pytorch [Paszke et al., 2019] for training neural networks.
-We used FrEIA [Ardizzone et al., 2018-2022] for constructing normalizing flows.
-Two-dimensional datasets
+Figure 8: Samples from the 5 two-dimensional datasets.
+B.1 Software Libraries
+We used the numpy (Harris et al., 2020) and scipy (Virtanen et al., 2020) Python libraries for basic numerical and graph operations and matplotlib (Hunter, 2007) for plotting.
+We used scikit-learn (Pedregosa et al., 2011) to compute nearest neighbor graphs and pytorch (Paszke et al., 2019) for training neural networks.
+We used FrEIA (Ardizzone et al., 2018-2022) for constructing normalizing flows.
+B.2 Two-Dimensional Datasets
 We use the following two-dimensional datasets.
@@
 The first two are explicitly Gaussian, the final three are GMMs with 50 components fitted using scikit-learn to data generated by adding noise to certain geometric structures.
-  • Standard normal
-  • Gaussian mixture with 3 components
-  • Circle
-  • 1 spiral
-  • 2 spirals
-Normalizing flow training
-We used the FrEIA library [Ardizzone et al., 2018-2022] to design normalizing flows using spline coupling blocks [Durkan et al., 2019]. We used the following hyperparameters:
+  1. 1.
+    Standard normal distribution
+  2. 2.
+    Gaussian mixture with 3 components:
+    • Means: $\mu_{1}=(2,1.4)$, $\mu_{2}=(6.5,6.3)$, $\mu_{3}=(8,1)$
+    • Covariances: $\Sigma_{1}=\begin{pmatrix}3&2.5\\ 2.5&3\end{pmatrix}$, $\Sigma_{2}=\begin{pmatrix}3&0\\ 0&3\end{pmatrix}$, $\Sigma_{3}=\begin{pmatrix}2&-0.8\\ -0.8&2\end{pmatrix}$
+    • Weights: $(0.25,0.5,0.25)$
+  3. 3.
+    Circle with radius 1 and Gaussian noise ($\sigma=0.08$)
+  4. 4.
+    Single spiral with 1.75 rotations ($3.5\pi$ radians) and Gaussian noise ($\sigma=0.05$)
+  5. 5.
+    Two interleaved spirals with 1 rotation each and Gaussian noise ($\sigma=0.045$)
+The spirals reach a radius of 1 after a single rotation.
+B.3 Normalizing Flow Training
+We used the FrEIA library (Ardizzone et al., 2018-2022) to design normalizing flows using spline coupling blocks (Durkan et al., 2019). We used the following hyperparameters:
   • Blocks: 12 for two-dimensional data, 5 for standard normal in $D$ dimensions
@@
   • Adam optimizer
-Score model training
-We use fully connected networks to predict the score, with 6 hiddens layer with the width depending on the dimension of the input. We use the following hyperparameters:
-  • Hidden layer width (as a function of input dimension $D$)
+B.4 Score Model Training
+We use fully connected networks to predict the score, with 6 hidden layers of varying width depending on the input dimension. The architecture consists of:
+  • Input layer → 6 hidden layers → Output layer (dimension matching input)
+  • Hidden layer widths:
     • $3\leq D\leq 5$: 128
@@
 In all cases, the chosen model checkpoint was that which achieved the lowest training loss.
-Compute resources
+B.5 Compute Resources
 For the two-dimensional datasets, we performed all experiments on a machine with a NVIDIA GeForce RTX 2070 GPU and 32 GB of RAM. It took approximately 24 hours to run all experiments.
@@
 Appendix C Additional Experimental Results
-Figure 6: Increasing the number of nearest neighbors does not significantly improve the convergence of the power-weighted graph method on two-dimensional datasets.
-C.1 Convergence of alternative graph edge weightings
-Figure 7: Alternative edge weightings based on the nearest neighbor density estimator do not converge.
-The same weightings based on the ground truth density do converge.
-Experiments performend on the GMM two-dimensional dataset.
+C.1 MNIST experiments
+Table 1: For each digit, we compute the mean and standard deviation of the log distances between the cluster mean and 20 randomly sampled points of that class (intra-class log distance), as well as between the cluster mean and 20 randomly sampled points of other classes (inter-class log distance).
+We also compute the difference, which gives an indication of how well-separated the digits are in terms of log distance.
+Digit Intra-Class Log Distance Inter-Class Log Distance Difference (Inter–Intra)
+0 2.41 ± 0.40 3.73 ± 0.27 1.32 ± 0.49
+1 1.69 ± 0.49 3.73 ± 0.31 2.04 ± 0.58
+2 3.56 ± 0.29 4.05 ± 0.19 0.49 ± 0.35
+3 3.03 ± 0.33 3.78 ± 0.24 0.75 ± 0.41
+4 3.04 ± 0.38 3.77 ± 0.36 0.73 ± 0.53
+5 3.43 ± 0.35 4.02 ± 0.18 0.59 ± 0.39
+6 2.89 ± 0.41 3.79 ± 0.31 0.91 ± 0.51
+7 2.86 ± 0.47 3.81 ± 0.30 0.95 ± 0.56
+8 3.29 ± 0.33 3.74 ± 0.21 0.46 ± 0.39
+9 2.76 ± 0.45 3.56 ± 0.38 0.80 ± 0.59
+Figure 9: Each row visualizes an MNIST geodesic, obtained by relaxation (Algorithm 1) using a learned score model within a 10-dimensional autoencoder latent space.
+We train an autoencoder to compress MNIST (LeCun et al., 2002) digits to a 10-dimensional space and subsequently train a sliced score matching model (Song et al., 2020) and normalizing flow (Dinh et al., 2017) on the latent distribution.
+Both encoder and decoder use fully connected layers with ReLU activations, with a single hidden layer of width 400.
+The score model has input and output dimension 10 and has hidden layers of width 128, 256, and 128, with SiLU activations.
+The normalizing flow is built with the FrEIA framework (Ardizzone et al., 2018-2022), made up of 4 affine coupling blocks, with a permutation between each block.
+Each block uses a coupling network with 2 hidden layers of width 128 and SiLU activations.
+All models are trained for 400 epochs with the Adam optimizer, the autoencoder with learning rate $10^{-3}$ and the score model and normalizing flow with learning rate $10^{-4}$.
+To compute distances between two points sampled from the latent distribution, we the same procedure as in the main text: i) construct a density-weighted graph (Algorithm 2) from the normalizing flow model ii) use Dijkstra’s algorithm to find the shortest path between the points according to the graph iii) solve the relaxation problem (Algorithm 1) with the score model, using the graph-based path as initialization iv) evaluate the length of the relaxed path using the normalizing flow model, according to Equation 2.
+In Table 1 we report mean and standard deviation of log distances between each digit cluster mean and 20 random samples of the same or other digits.
+We see that same-class distances are always lower than other-class distances, up to a safe margin of error.
+We also make the following observations: 1s cluster tightly, likely due to having fewer active pixels and therefore having higher likelihoods; 8s are more ambiguous, likely due to easy deformations into other digits (it is quite easy to remove some pixels to turn an 8 into a 3, 5, 6 or 9).
+We leave deeper analysis of Fermat distances on MNIST to future work.
+In Figure 9 we visualize geodesics between 5 randomly sampled pairs from the latent distribution.
+The geodesics are solved in the latent space and the resulting latent points are decoded to image space for visualization.
+We see that the geodesics follow high-density regions of the distribution, leading to interpretable and reasonable transitions between the digits.
+C.2 Convergence of Alternative Graph Edge Weightings
+Figure 10: Alternative edge weightings based on the nearest neighbor density estimator converge poorly.
+The same weightings based on the ground truth density converge well.
+Experiments performed on the GMM two-dimensional dataset.
 We try out other edge weights, as described in the main text.
@@
     Min: $\tilde{p}((X_{l}+X_{m})/2)=\min(\hat{p}(X_{l}),\hat{p}(X_{m}))$
-We find that all 4 have very similar performance to the power-weighted graph, whereas the equivalent estimators using the ground truth density quickly converge. See Fig. 7.
-Figure 8: Detailed plot of convergence behavior in higher dimensions, unscaled metric.
-We see that all graph-based methods plateau in performance, whereas the relaxation-based method continues to show good convergence.
-Figure 9: Detailed plot of convergence behavior in higher dimensions, scaled metric.
-We see that all graph-based methods plateau in performance, whereas the relaxation-based method continues to show good convergence.
+We find that all 4 have very similar performance to the power-weighted graph, whereas the equivalent estimators using the ground truth density quickly converge. See Figure 10.
+C.3 Trade-off Between Log Density and Score Estimation
+Figure 11: Mean integrated squared error (MISE) for log density and score estimation using KDE with varying bandwidth.
+The optimal bandwidth for log density estimation is smaller than the optimal bandwidth for score estimation.
+We speculate that there is a fundamental trade-off between accurately estimating log densities and their gradients (score functions) that applies to all density estimation methods, including the normalizing flows used in our main experiments.
+To illustrate this principle in a simple setting where we can compute exact errors, we analyze kernel density estimation (KDE) of the standard normal distribution in one dimension using Gaussian kernels, with 1000 samples drawn from the distribution.
+While this example is much simpler than the high-dimensional problems and neural network models considered in the main text, it clearly demonstrates how optimizing for density estimation can lead to poor gradient estimates and vice versa.
+For a set of samples $\{x_{i}\}_{i=1}^{n}$, the KDE estimate with bandwidth $\sigma$ is:
+$\hat{p}(x)=\frac{1}{n\sigma\sqrt{2\pi}}\sum_{i=1}^{n}\exp\left(-\frac{(x-x_{i})^{2}}{2\sigma^{2}}\right)$ (34)
+The log density and score estimates are then:
+$\log\hat{p}(x)$ $=-\log(n\sigma\sqrt{2\pi})+\log\sum_{i=1}^{n}\exp\left(-\frac{(x-x_{i})^{2}}{2\sigma^{2}}\right)$ (35)
+$\hat{s}(x)$ $=-\frac{\sum_{i=1}^{n}(x-x_{i})\exp\left(-\frac{(x-x_{i})^{2}}{2\sigma^{2}}\right)}{\sigma^{2}\sum_{i=1}^{n}\exp\left(-\frac{(x-x_{i})^{2}}{2\sigma^{2}}\right)}$ (36)
+Figure 12: Comparison of KDE estimates using different bandwidths.
+Top row: Log density estimates.
+Bottom row: Score estimates.
+Left column: Using the bandwidth that minimizes log density MISE.
+Right column: Using the bandwidth that minimizes score MISE.
+Ground truth shown in dotted orange, KDE estimate in solid blue.
+The optimal bandwidth for log density leads to noisy score estimates, while the optimal bandwidth for score estimation leads to oversmoothed log densities.
+As shown in Figure 11, the optimal bandwidth for minimizing the mean integrated squared error (MISE) of the log density estimate is smaller than the optimal bandwidth for minimizing the MISE of the score estimate.
+This reflects a fundamental trade-off: accurate density estimation requires capturing fine details of the distribution, while accurate score estimation requires smoothing out noise to obtain reliable gradients.
+Figure 12 visualizes this trade-off.
+When using the bandwidth optimal for log density estimation, the score estimates are quite noisy.
+Conversely, when using the bandwidth optimal for score estimation, the log density estimates are oversmoothed.
+This observation helps explain why we obtain better results using dedicated score models rather than differentiating density models.